### v1

# Malsight: Exploring Malicious Source Code and Benign Pseudocode for Iterative Binary Malware Summarization

Abstract
Binary malware summarization aims to automatically generate human-readable descriptions of malware behaviors from executable files, facilitating tasks like malware cracking and detection. Previous methods based on Large Language Models (LLMs) have shown great promise. However, they still face significant issues, including poor usability, inaccurate explanations, and incomplete summaries, primarily due to the obscure pseudocode structure and the lack of malware training summaries. Further, calling relationships between functions, which involve the rich interactions within a binary malware, remain largely underexplored.
To this end, we propose Malsight, a novel code summarization framework that can iteratively generate descriptions of binary malware by exploring malicious source code and benign pseudocode. Specifically, we construct the first malware summaries, MalS and MalP, using an LLM and manually refine this dataset with human effort. At the training stage, we tune our proposed MalT5, a novel LLM-based code model, on the MalS dataset and a benign pseudocode dataset. Then, at the test stage, we iteratively feed the pseudocode functions into MalT5 to obtain the summary. Such a procedure facilitates the understanding of pseudocode structure and captures the intricate interactions between functions, thereby benefiting the usability, accuracy, and completeness of summaries. Additionally, we propose a novel evaluation benchmark, BLEURT-sum, to measure the quality of summaries. Experiments on three datasets show the effectiveness of the proposed Malsight. Notably, our proposed MalT5, with only 0.77B parameters, delivers comparable performance to much larger ChatGPT3.5.
Index Terms:
Malware, Code Summarization, Binary Code
1 Introduction
The AV-TEST Institute [1] recently reported that over 450,000 new malicious files and potentially unwanted applications are registered daily, showing a high demand for malware understanding. Binary malware summarization [2] is a reverse engineering [3] task that aims to automatically generate concise human-readable descriptions of binary executable malicious files. The summarization provides security analysts with a quick understanding of the malware’s functionality and patterns when source code is unavailable, thereby benefiting a wide range of applications such as malware cracking [4] [5] [6], malware family classification [7] [8], binary code similarity detection [9] [10] [11], and large-scale malware behavior analysis [12] [13] [14].
Figure 1: The comparison of source code (left) and its pseudocode (right). The pseudocode includes significantly more content and a more complex structure, and it also strips key semantic cues such as function names.
Existing reverse engineering tools, such as IDA [15] and Ghidra  [16], can decompile executables into higher-level C-like pseudocode, while they still lack easy-to-understand semantics information. Consequently, a line of efforts attempts to generate human-readable summaries based on pseudocode. Early studies rely on manual parsing or rule-based summary generation [17] [18]. Recent large language models (LLMs), such as BinT5 [19], HexT5 [20], CodeGen [21], and WizardCoder [22], have shown great potential to produce more informative summaries. However, these data-driven approaches still face critical issues, including poor usability, inaccurate explanations, and inaccurate explanations [2]. Figure 1 shows the underlying reasons for the above issues by comparing the source code of the function “initLevel” to the corresponding pseudocode. We observe that the pseudocode presents 1) significantly more content that increases from 20 lines in source to 117 lines in the pseudocode, 2) a more complex and obscure structure with multi-level nesting and entangled logic. The pseudocode involves 29 more calls and 29 more if statements compared to the source code at the left, 3) stripping key semantic cues such as variable names and function names. For example, the function “initLevel” in source code is transferred to a meaningless symbol “sub_404018”.
To address the above challenges, we present Malsight, a novel binary malware summarization framework that can iteratively generate descriptions of executable malware by exploring malicious source code and benign pseudocode. The proposed Malsight involves three key ingredients, including a malware dataset MalS, an LLM-based malware summarization model MalT5, and an evaluation metric BLEURT-sum. We describe the workflow of the proposed framework in four steps as follows.
Constructing MalS: As an LLM-based summarization model heavily relies on high-quality annotations to learn to align with domain-specific knowledge, it necessitates high-quality malware pseudocode summaries to fine-tune the LLM. However, the public malware pseudocode summarization dataset is unavailable so far, and building such a benchmark is quite challenging as it requires huge human involvement for accurate annotations. Figure 1 illustrates three challenges of understanding malware pseudocode. To tackle this issue, we alternatively construct MalS, a large-scale summarization dataset using an LLM model, and malicious C language source code crawled from GitHub. The proposed MalS involves nearly 90,000 malware source functions, with 20 types of malware functions. We also construct a small dataset MalP for testing. We detail such a procedure in Section 4.3.
Training MalT5: We use CodeT5+ [23] as the foundation model of our MalT5. We sequentially fine-tune the proposed MalT5 model on the MalS dataset and an existing benign pseudocode summarization dataset [19]. The underlying intuition is that the malicious semantic knowledge from malware source code summarization and function patterns from benign pseudocode summarization, which are learned from the above two datasets, respectively, can be transferred to the generation of malware pseudocode. By doing so, we can properly mitigate the issue of unavailable malware pseudocode summarization datasets. More details are available in Section 4.4.
Performing Generation: We use an existing tool [15] to generate pseudocode of a binary file and then generate summaries using the MalT5 model. We first use IDA to construct the malware call graph and then develop an algorithm to transform the graph into a function list in reverse order. Then we iteratively fed the first function in the list to MALT5 to generate the summary. More details are provided in Section 4.1 and 4.2.
Conducting Evaluation: Previous work [24] indicated that existing metrics for generation tasks, such as Bilingual Evaluation Understudy (BLEU) [25], Metric for Evaluation of Translation with Explicit ORdering (METEOR) [26], Recall-Oriented Understudy for Gisting Evaluation-Longest Common Subsequence (ROUGE-L) [27], may not well-fit for evaluation of the binary malware summarization. We thus employ BLEURT-sum, which is more sensitive to the quality of the pseudocode summary, thereby benefiting the evaluation in practice. More descriptions are given in Section 5.1.
We conduct experiments on three datasets to verify the effectiveness of the proposed Malsight framework for binary malware summarization. The contribution of this paper can be summarized as follows [^1] We will release our Malsight to contribute to the community..
  • A binary malware summarization Framework. We propose Malsight, a novel framework that can iteratively generate descriptions of binary malware by exploring malicious source code and benign pseudocode. Our MalT5 can tackle the challenges of entangled logic and stripped semantics in pseudocode.
  • Large-scale datasets for binary malware summarization. We propose MalS and MalP, two novel datasets that can be used for the LLM training and testing of an LLM of binary malware summarization. To the best of our knowledge, the two datasets are the first in the field, involving nearly 90,000 malicious source functions and 20 types. Our MalS and MalP can serve as a benchmark for various binary malware understanding tasks.
  • An LLM-based binary malware summarization model. We propose MalT5, a novel LLM for the summarization task. The proposed MalT5 is lightweight, with only 0.7B parameters.
  • An evaluation metric for the task: We present BLEURT-sum, a novel evaluation metric that is more sensitive to the quality of pseudocode summarization.
  • Extensive experiments. We conduct extensive experiments on three datasets and provide case studies to show why the proposed framework performs best among all baselines. Results show that our MalT5 achieves comparable performance to ChatGPT3.5.
2 Background
2.1 Malware Analysis Engineering
The field of malware analysis engineering focuses on analyzing the functionality of malware by examining its binaries, typically through static analysis methods that involve observing assembly code or pseudocode [28].
2.1.1 Binary Decompilation
Decompilation [29] converts executable files into human-readable pseudocode [30], which is more concise and structured than disassembled assembly code. Unlike disassembly, which maps instruction encoding directly to assembly statements, decompilation relies on algorithms and patterns (e.g. R2 [31], IDA [15], Ghidra [16]) and emerging methods using LLMs [32]. However, pseudocode lacks semantic information such as function names. Decompiled function names are often unreadable (e.g., sub_4061C0 in IDA Pro) [33], providing a little useful pieces of information for further analysis.
2.1.2 Human Static Analysis
In static analysis, human experts start analyzing from the function entry point [34], inferring functionality from system Application Programming Interface (API) calls, string information, and pseudocode logic. Their main challenge is accurately identifying the core function [35] among numerous functions and methodically tracing the function call [36] process to understand the functionality comprehensively. To assist in this process, we developed Machine Learning-based (ML-based) Malsight, which optimizes and facilitates binary malware analysis.
2.2 NLP Technologies
In the Malsight process, we use Bidirectional Encoder Representation from Transformers (BERT) [37] and Text-to-Text Transfer Transformer (T5) [38] architecture language models to complete specific tasks. For the core code summary task, we build a CodeT5+ model combined with transfer learning.
2.2.1 BERT Family
BERT is a large-scale transformer-based language model pre-trained on a wide corpus of text using a self-supervised learning approach. The design of Masked Language Modeling (MLM) and Next Sentence Prediction (NSP) training tasks makes BERT perform well in the tasks of Sequence Labeling (SL), such as Named Entity Recognition (NER).
Building on BERT, CodeBERT [39] learns code semantics through Code-Conditioned masked language modeling (CMLM) and natural language documentation generation (NLG). CodeBERT has been shown to perform well on code-related tasks, and since it is derived from BERT, we have reason to believe that this model can be fine-tuned to solve the problem of SL in pseudocode.
2.2.2 T5 Family
T5 [38], or Text-to-Text Transfer Transformer, is a sequence-to-sequence model based on the Transformer architecture that unifies various Natural Language Processing (NLP) tasks into a single framework, including text classification, question answering, summarization, translation, and text generation.
CodeT5 [40] is an encoder-decoder model supporting code understanding and generation, built on the T5 architecture. It uses Natural Language-Programming Language (NL-PL) bimodal data for pre-training with identifier tagging and masked identifier prediction tasks. CodeT5+ [23] introduces greater architectural flexibility and additional pre-training tasks, with instruction tuning to enhance alignment with natural language instructions. This results in significant performance improvements on various code-related tasks.
Previous works like HexT5 [20] and BinT5 [19] developed datasets to train models for binary code understanding, including code summarization tasks. These efforts demonstrate the potential of T5-based models in binary code summarization for malware analysis.
2.2.3 Transfer Learning
Transfer learning involves training a model on a source task with abundant labeled data to learn general features. When it is difficult to obtain sufficient datasets for training, transfer learning can be used to supplement them with similar or related datasets [41].
In Malsight, we fine-tune the CodeT5+ [23] model to achieve transfer learning from the source code summarization task to the decompiled code summarization task. Besides, we use dynamic and static annotation to implement feature enhancement to compensate for the poor transfer effect caused by the highly limited similarity between the source code and the stripped decompiled code.
2.3 Code Summary Evaluation
In code summary model evaluation, NLP text similarity algorithms compare generated results with a reference test set, replacing costly human evaluations. These algorithms are categorized into word overlap and word embedding measures.
2.3.1 Words’ Overlap Measure
Early text similarity measures like BLEU [25] and ROUGE [27] rely on word n-gram overlap between generated and reference text, with BLEU focusing on precision and ROUGE on recall. However, they lack semantic understanding. METEOR [26] integrates n-gram overlap and semantic similarity using WordNet, providing additional semantic insight.
Recent work [42] highlights limitations of words’ overlap in code summary tasks. It shows that similar structures may yield high similarity scores despite differing semantics.
2.3.2 Words’ Embedding Measure
The words’ embedding measure evaluates semantic similarity by analyzing the distance between sentence embeddings in a vector space, often utilizing neural network learning.
word2vec [43] is a static embedding model that represents words as points in a vector space, facilitating the proximity of semantically similar words. MoverScore [44] uses an n-gram optimized Word Mover’s Distance (WMD) [45] to measure similarity and employs various embedding models like ELMo [46] and BERT.
BLEURT [47] stands out as a versatile metric designed for assessing various natural language generation tasks, which combines the advantages of both Words’ Overlap Measure and Words’ Embedding Measure. It achieves this by integrating diverse lexical and semantic-level supervision signals into its pre-training process and leveraging synthetic data based on pre-trained BERT, ensuring its effectiveness and versatility in various evaluation scenarios.
Figure 2: Workflow of Malsight. The procedure involves three steps: graph traversal, annotation generation, and model summary.
3 Motivation and Overview
The construction of the code summary framework mainly includes annotation generation and code summary model construction, as shown in Figure 2. During the evaluation phase, we tested several evaluation methods and found a reasonable way to build an evaluation model for the code summary task. Simultaneously, our work involves the construction of multiple datasets (for subsequent stages of training and evaluation of transfer learning-based models).
3.1 Code Summarization Process
The code summary task is split into three steps, which are function list extraction, annotation generation, and code LLM summary.
3.1.1 Function List Extraction
As mentioned, existing code summary methods for binary focus only on the internal information of the function. We introduced the call relationship between functions and worked on the entire binary as the processing unit. In other words, when function func_E in Figure 2 calls func_F, it is difficult for the subsequent code summarization model to correctly summarize the functionality of func_E without any information about func_F. (We assume that the function name of func_F has been corrupted.) Constructing a list of reverse call sequential relational functions provides a basis for the subsequent recovery of sub-functions functionality.
3.1.2 Annotation Generation
Iterate through the list of functions (assuming Fun_F has been processed), and Fun_E will first be added with annotations by the static annotator and the dynamic annotator, respectively. Fun_E uses the static annotator to obtain static annotations based on the internal information of the function code, while the dynamic annotator adds dynamic annotations based on the generated summary of the sub-function (Func_F) in the function. The program, in other words, sequentially restores functions according to the Control Flow Graph (CFG) diagram from the outermost to the innermost and passes function summary results inward.
3.1.3 Code LLM Summary
Fun_E(annotated) is then fed into the code summary model for final code summary generation. In our work, we use transfer learning to adapt the model to both the functionality of malware functions and the structural features of decompiled pseudocode. Based on the CodeT5+ model, we have fine-tuned the code summary task. The tokenizer splits the code into tokens and embeddings, incorporating a self-attention mechanism into a complete vector in the encoder. The decoder outputs a fine-tuned prediction based on the code summary.
Figure 3: A Sample Of Existing Evaluation Results. Two unrelated pairs of two sentences form Sample I, and two related pairs form Sample II, but the evaluation results contradict expectations.
3.2 Evaluation Method
Our research has found that existing methods can not simultaneously measure the meaning, structure, word frequency, and other features of the reference sentence and the candidate sentence, so the model’s performance may be misjudged. Taking Figure 3 as an example, two examples show the evaluation results of BLEU, Meteor, and ROUGE-L on two pairs of real code summaries. The figure shows that two code summaries without any semantically related results in high evaluation scores (blue-framed), while two semantically similar code abstracts receive low scores (red-framed), demonstrating the shortage of existing methods. In the following work, we construct an ML-based code summary evaluation method BLEURT-sum by constructing a set of positive and negative samples composed of related sentence pairs and unrelated sentence pairs. We evaluated the usability of the model and prior art by measuring their ability to distinguish between positive and negative samples.
3.3 Datasets Construction
For the two core tasks mentioned above, binary code summary and code summary model evaluation, we build corresponding datasets.
3.3.1 Dataset For Code Summary Model
In order to avoid the data shift problem, the training of the code summary model requires a large dataset of malware pseudocode summary. Unfortunately, malware datasets are typically represented as collections of compiled binary files [48], with the binary code stripped, and possibly structurally confused. Consequently, the creation of code summary datasets for binary malware could be unfeasible without resorting to labor-intensive manual summarization.
In this paper, our key insight is that the code summary model requires two capabilities, understanding of Malware Functionality and adaptability to disassembled pseudocode formats (including the ability to deal with annotated code). Therefore, We build source-based malware datasets and pseudo-code-based benign software datasets to train the model on these two capabilities separately. Based on Sourcefinder [49], we were able to find malware source repositories from GitHub. We generate descriptive labels for the extracted malware functions using a sophisticated language model, followed by manual verification and optimization. Meanwhile, we use the Capybara dataset provided by BinT5 [19] (a benign software dataset) to train the model adaptability to pseudocode structure.
3.3.2 Dataset For Evaluation Model
In the evaluation phase, the evaluation method is used to measure the similarity between the model generation results and the reference results to evaluate the quality of the model generation. Given two sentences (generation results and the reference results) $S_{g}$ and $S_{r}$, most evaluation methods output a score $Score$ as the evaluation result. Therefore, if considering the use of machine learning methods, it is necessary to construct a dataset in $\{S_{g},S_{r},Score\}({Score}\in[0,1])$ format. The challenge is that when a dataset of $\{S_{g},S_{r}\}$ is obtained, it is a difficult job to obtain an accurate $Score$. In our subsequent work, we propose a reasonable algorithmic flow for constructing labeled datasets $EvaS$.
3.3.3 Dataset For Static Annotater
In annotation generation process, the static annotater includes a core information extraction module (described in detail in Section 4.2.1). Due to the difficulty in accurately completing the required functions using static methods, we use a machine learning model to complete the sequence labeling task of the pseudocode. By constructing the dataset AnnoS, we have constructed the dataset required for model training and testing.
TABLE I: The Proposed Datasets
Sets for code summary model
Datasets Size(functions) Code language Annotated? Usage
MalS 89,609 C No Train phase1
MalP 500 pseudo Yes Test
BenignC 96,835 pseudo Yes Train phase2
Sets for annotation extractor model
Dataset Size(functions) Code language Anno num(avg) Usage
AnnoS 95,000 pseudo 3.87 Train & Test
Sets for evaluation model
[…]
EvaS 127,510 1:1 9.6 Train & Test
To sum up, we mainly completed the construction of three sets of datasets in different application fields, as shown in Table I.
4 Code Summarization Workflow
Following our breakdown of the malware code summary task in Figure 2, our implementation first extracts the reverse function list, and then sequentially generates static and dynamic annotations for the items in the function list, and finally passes them into the code summary model.
In this process, we completed the training and designing of two Domain-Specific large model, the applying of a General-Purpose large model and the implementation of several algorithms. we cover the implementation of each step separately in this section.
4.1 Function List Extraction
As mentioned earlier in Section 3.1, in the first step of the workflow, we extract the list of reverse functions from the CFG of the malware binary.
Since the existing methods [50] do not give a completely accurate CFG extraction flow, we implement a pluggable CFG extraction module. It is used to provide us with a processing scheme from binary file $B_{Mal}$ to digraph $G_{M}$ as the CFG. By an inverse topological traversal algorithm, it is extracted from CFG in the opposite direction of the call chain, expressed as $L_{G_{M}}=[f_{1},f_{2},...,f_{n}](n=G_{M}.vertices)$, Where $f_{i}$ represents the $i$ th function of inverse topological order of $G_{M}$.
In this study, algorithm $REsort$ was constructed, expressed as $L_{G_{M}}=REsort(G_{M})$. (See Appendix B for details) By applying algorithms such as Tarjan [51], Dijkstra [52], and Depth First Search (DFS), this approach successfully addressed the obstacles caused by cyclic calls and partially connected graphs in the process of generating reverse order lists. Function columns $L_{G_{M}}$follow the order from outside to inside in the function call diagram to ensure that the later-called function is first in the function list and is processed first in subsequent steps.
The following process takes the function instances from $L_{G_{M}}$ in a forward order to achieve the order of restoration from the outer layer of the CFG diagram to the inner layer, specifically, recovering from the outer API call to the main function.
4.2 Annotation Generation
Figure 4: Annotation Workflow. If the function currently being processed is Func, its dynamic annotation comes from the child function it calls callee, and the static annotation is generated in three steps. Func’s code summary is then passed to the Caller as a dynamic annotation.
In the order of traversal provided by the reverse function list, annotations are added to each function in turn to provide richer information. Annotations can be divided into static and dynamic types.
We have observed that attackers frequently employ techniques like stripping to hinder reverse engineering efforts. This process removes crucial symbolic information, such as identifier names, from the binaries, leading to significant semantic loss in the generated pseudocode. As a result, code summary models for malware face substantial challenges, and the performance of code summarization tasks is adversely affected.
To address this issue, we propose utilizing dynamic and static annotation to supplement the semantic information and enhance the features of the stripped pseudocode as Figure 4 shows. This approach aims to compensate for the poor transferability caused by the stark dissimilarity between the source code and the stripped decompiled code.
4.2.1 Static Annotation
Based on our long-term exploration of pseudocode for stripped malware, we deem that although the stripped pseudocode has a serious semantic loss, some extremely critical API calls (such as operating system APIs) and some special forms of strings that are preserved after stripping provide us with ideas for behavior analysis and semantic recovery of malware.
Consequently, we consider building a static annotation module to provide additional information for subsequent code summaries. In general, the static annotation process can be divided into three parts: sequence labeling, online retrieval and annotation generation.
Sequence labeling model: By manually labeling approximately 300,000 tokens within nearly 80,000 functions, we construct a labeled dataset $Set_{SL}$ based on $BenignS$ to train the sequence labeling model.
Formally, the function $f_{i}$ is first sliced into an n-token code sequence by the tokenizer $T$. The n-token code sequence $s_{i}=T(f_{i})=\{t_{0},t_{1},...,t_{n-1}\}$ and the labels of the tokens $L_{i}=\{l_{0},l_{1},...,l_{n-1}\}$ are combined into $d_{i}=\{(t,l)|t\in s_{i},l\in L_{i}\}$. Equation (1) formalizes this dataset.
$Set_{CSL}=\{d_{0},d_{1},...,d_{N}\}$ (1)
Subsequently, we opted to utilize the CodeBERT [39] model for training the sequence labeling task using this dataset CSL.
$B$ is the CodeBERT base model, and $C$ is the classifier head. As a sequence is inputted, the complete model outputs the predicted labels of its tokens. Equation (2) and (3) formalize this.
$B(s_{i})=\{o_{0},o_{1},...,o_{n-1}\}\vspace{-0.2cm}$ (2)
$C(o_{i})=y_{i}$ (3)
We further formalize the target function in equation (4), where $l_{i_{c}}$ is the truth label and $y_{i_{c}}$ is the Softmax probability for the $c^{th}$ class.
$LF=-\sum\limits_{c=0}^{2}l_{i_{c}}\log{y_{i_{c}}}$ (4)
By optimizing $LF$, $B$ and $C$ are trained simultaneously. This necessitates that the model effectively classifies code-tokens to accurately label the key API calls and special strings within the pseudocode.
Label to annotation: Once we can get the key API calls and special strings in the stripped pseudocode, we use GitHub Code Search [53] to retrieve the relevant context in the GitHub repositories. In the implementation, we keep the first three blocks of the search results (this is because the code search has already sorted the relevance of the results [54]).
The outcome of random sampling and manual discrimination reveals that approximately 54.8% of the function contexts contain code comments closely associated with the functionality of the function. For the remaining functions, nearly 90% also offer contextual information related to the function’s operation, such as parameter names, interconnected functions, and processing logic. Only a small number of functions yield invalid search results.
The filtered and preprocessed code snippets will be continuously input into the prompt-based generic model for generating static annotation.
4.2.2 Dynamic Annotation
As shown in Figure 4 (the blue parts represent the steps in which the annotation was added), the summary of the callee is provided to the caller as a complement to the semantic information, which we define as dynamic annotation. This is consistent with the actual analysis flow of binary malware analysis by reverse workers, i.e., analyzing the call relationship from the inner layer of the CFG diagram to the outer layer (corresponding function list generated in Section 4.1) and summarizing the function from the outside in (corresponding to the passing of dynamic annotation). In this way, we can make full use of the dynamic behavior characteristics implied by the call relationships between functions in the pseudocode.
4.3 Building Malware Datasets
In the traditional scheme of building datasets for decompiled code, the datasets are built at the function level [55]. The source function $f_{n}^{Source}$ is compiled and linked with other modules to generate an executable file, as $f_{n}^{Bin}$, and then decompiled to obtain the pseudocode form $f_{n}^{pseudo}$ of the corresponding function. Equation (5) formalizes this dataset, where $SUM()$ is the extraction method for the code summary.
$Set_{ideal}=\{(f_{n}^{pse},f_{n}^{sum})|f_{n}^{sum}=SUM(f_{n}^{pse})\}$ (5)
Out of 2,289 GitHub repositories that were determined to be malware, we extracted close to 30K functions. We filter for functions that repeat, shorter than five lines, and format-challenged functions, resulting in a dataset of 89,609 functions. (The lack of strict filtering may lead to overlap between the train and test sets, consequently yielding inflated results.)
Unlike benign open-source projects that are well maintained, most of the functions we extract do not have comments in context for us to label as code summaries. Fortunately, the semantic information in the source code is rich enough that we used a well-designed prompt to complete the code summary for us via GPT3.5-Turbo [56]. In this way, we extract dataset MalS in equation (6) format. Further, we build a 500-function dataset MalP for testing our model. The reason why dataset MalP has a relatively small quantity is that it was obtained by manually compiling and decompiling the git repository. MalP is compiled from the makefile provided by the developer, so we have not mentioned the configuration related to compilation optimization.
$Set_{MalS}=\{(f_{n}^{Sou},f_{n}^{sum})|f_{n}^{sum}=SUM(f_{n}^{Sou})\}$ (6)
$Set_{MalP}=\{(f_{n}^{pse},f_{n}^{sum})|f_{n}^{sum}=SUM(f_{n}^{Sou})\}$ (7)
As mentioned above, the malware training set we built was made up of source code, so another dataset was needed to help our model understand the stripped function features. In the transfer learning option, we used the Capybara dataset (a benign software dataset) provided by BinT5 [19] and annotated it to provide our model with adaptations to the annotated code summary task. It is provided as $Set_{Capybara}$, and we process it as $Set_{benignC}$, where the $ANN()$ is the static annotation generation method to form PL-NL bimodal data $f_{n}^{ann}$.
$Set_{Capybara}=\{(f_{n}^{pse},f_{n}^{sum})|f_{n}^{sum}=SUM(f_{n}^{Sou})\}$ (8)
$Set_{benignC}=\{(f_{n}^{ann},f_{n}^{sum})|f_{n}^{ann}=ANN(f_{n}^{pse})\}$ (9)
In the process of training the code summary model, we use $Set_{MalS}$ and $Set_{benignC}$ to complete the transfer learning process, and $Set_{MalP}$ to test during the evaluation phase.
Figure 5: Code Summary Topic Distribution. The results of topic analysis for ${Set_{MalS}}$ indicate that its distribution conforms to the functional distribution of malware functions.
For $Set_{MalS}$, 25% of the samples were extracted for thematic analysis of function functionality, as shown in Figure 5. Our analysis results confirm that $Set_{MalS}$ and $Set_{BenignC}$ have significantly different theme distributions. Among a large number of security-related functions in the former, there are an unignorable number of codes for lock, permissions, and encryption, while in the latter, there are a large number of codes related to game logic and driver calls which irrelevant to malware.
4.4 Code Summary Model
In Malsight, we fine-tune the CodeT5+ model to enable transfer learning from source code summarization to decompiled pseudocode summarization. To achieve this, we have designed two distinct phases of fine-tuning to facilitate smooth transfers that accommodate variations in data characteristics and distribution biases across different datasets.
Phase 1. In the initial phase, we fine-tune the model that uses the dataset we built consisting of malware source code $Set_{MalS}$. The source code of malware has a similar semantic structure to the code in ordinary scenarios, but it also has behavioral or semantic features that the latter does not exist, which are represented by some code fragments with malicious purposes, such as self-replication and propagation, illegal access to system resources, and vulnerability exploitation. Through this phase of fine-tuning, the model can learn the behavioral or semantic features of these malicious codes. In this phase, the poor effects of data shift are addressed by using $Set_{MalS}$ with the same functional distribution as malware in the real world.
Phase 2. In this phase, we ask the model to learn the corresponding semantic information from the pseudo-code and the annotation text simultaneously to better assist the model in generating high-quality code summaries. We use the PL-NL bimodal dataset BenignC to fine-tune the encoder of the model which we fine-tuned in the previous phase in case the decoder’s parameters are frozen. By only fine-tuning the encoder, we allow the model to adjust itself to better suit the specific task at hand, without updating the weights too much. This can not only lead to a more robust and generalized understanding of the data but also shorten the training time due to the reduction in parameters. Especially, the raw input is processed into a standardized format as equation (10) where $t_{c_{i}}$ is the tokens from a code sequence, $t_{a_{i}}$ is the tokens from the corresponding annotation, $t_{sep}$ is a special token in CodeT5+ to separate the inputs in different modes. Inputs in such a format can assist the encoder in distinguishing the difference between two modes and learning the correlation between them.
$f_{n}^{ann}\to\{t_{c_{0}},...,t_{c_{n-1}},t_{sep},t_{a_{0}},...,t_{a_{m-1}}\}$ (10)
After the above two phases of fine-tuning, the model basically has the ability to accept bimodal input composed of pseudocode and annotation text and generate code summary. We use this model to summarize the entire malware in reverse of the function call order.
5 Evaluation Method
As mentioned in Section 3.3.2, the evaluation algorithm for code summary tasks should accept the reference as input and separate available and unavailable generated results. In this section, we introduce our exploration of code summary dataset construction (EvaS) and evaluation model construction (BLEURT-sum) respectively.
5.1 Evaluation Dataset Construction
Utilizing the code summary from the MalS dataset as a foundation for our research, we curated a positive and negative sample pair for the tuning of our new evaluation model BLEURT-sum. The positive sample consisted of two code summary result sentences sharing the same meaning and was initialized as $\{S_{g},S_{r},1\}$, while the negative sample comprised two randomly different sentences and was initialized as $\{S_{g},S_{r},0\}$. $S_{g}$ and $S_{r}$ represent generated statement and reference statement, respectively.
In order to build the dataset in $\{S_{g},S_{r},Score\}({Score}\in[0,1])$ format, one possible idea is to build an algorithm, $Score=GenSim(S_{g},S_{r},0or1)$, to automatically generate the label $Score$ required for model training. Therefore, the key task is to construct a reliable $GenSim()$ function. When $S_{g}==SUM(f_{n})$, contrusted $Score$ can be expressed using equation (11), considering that the SUM() function produces different outputs when faced with the same input, which is a non-deterministic function.
$S_{r}\!=\!\text{SUM}(f_{n})\!\implies\!Score\!=\!\begin{cases}1\!&\!\text{if }S_{g}=\text{SUM}(f_{n})\\ 0\!&\!\text{if }S_{g}=\text{SUM}(\neg f_{n})\end{cases}$ (11)
Since when $\{S_{g},S_{r}\}$ pairs were built, no sentence structure dependencies are taken into account, which means the equation (11) of Score can contain a few words overlap-based features (§ 2.3), the natural consideration is to combine the characteristics of sentence structure and semantics. Our idea is to solve the proportion of semantic information and sentence structure information in sentence similarity evaluation.
Taking the original 0,1 tag as the semantic feature, we further extract the static feature to get a Score, multiply the two by the corresponding proportion respectively, and then add them to get $Score$. Assuming that the proportion of semantic information is $p$, the structural feature calculation function is $Struc()$, the following equations formalize this Score construction method.
$s_{f}=Struc(S_{g},S_{g})$ (12)
$Score\!=\!\begin{cases}p+(1-p)*s_{f}\!&\!\text{if}S_{g}=\!\text{SUM}(f_{n})\\ (1-p)*s_{f}\!&\!\text{if}S_{g}=\!\text{SUM}(f_{\neg n})\end{cases}$ (13)
In the implementation, we utilize the arithmetic average of BLEU, ROUGE-L, and METEOR as the metric for $Struc()$ (they have been normalized to the zero-one interval using a uniform probability distribution, respectively). Then we solve the minimum square error and get the semantic and structural ratio close to 1:4, which makes $p=1/5$.
5.2 Evaluation Model
Based on our insights into the evaluation model, it should be able to combine both structural and semantic features and give an evaluation score for the specific task of code summary.
BLEURT is based on BERT and adds additional pre-training steps between pre-training and fine-tuning to the synthesized data. Synthetic data is generated by perturbing sentence pairs $<z,\widetilde{z}>$, where $z$ and $\widetilde{z}$ are randomly selected sentence pairs. In the additional pre-training of BLEURT, a series of pre-training signals $(\tau_{1},\tau_{2},...\tau_{9})$ to align the model with the desired result. BLEURT uses the sentence pair scores of BLEU, ROUGE, and BERTScore as signals $\tau_{1}$ to $\tau_{3}$, and uses the back translation processing sentence pairs to generate $\tau_{4}$ to $\tau_{7}$.
The rich training signals ensure the universality of BLEURT and the comprehensiveness of the evaluation angle. We then further fine-tune the BLEURT model on the code summary sentence pair dataset. First, the generated text and the reference text are input together into the model and the vector is generated as shown in equation (14). This solution is called BLEURT-sum.
$v_{[CLS]},v_{S_{g1}},...,v_{S_{gn}},...,v_{S_{rn}}=BLEURT(S_{g},S_{r})$ (14)
Further, the model uses the CLS vector to add antecedents to obtain the scores predicted by the model, as shown in equation (15).
$\hat{Score}=f(S_{g},S_{r})=W\tilde{v}_{[CLS]}+b$ (15)
Finally, the model starts to complete the loss calculation based on the loss and then carries out the gradient descent (equation (16)).
$loss=\frac{1}{N}\sum\limits_{n=1}^{N}\left|\left|Score-\hat{Score}\right|\right|^{2}$ (16)
In existing works, Word overlap measures are commonly used in text similarity evaluation but perform poorly for code summary evaluation. Code summaries often include keywords like “retur” and “initializ”, which do not significantly contribute to the overall meaning. These keywords can inflate overlap rates and lead to misleadingly high scores. Therefore, BLEURT-sum on the one hand still considers Word overlap measures, and on the other hand introduces more dimensions, which greatly improves the performance in code summary evaluation tasks.
6 Experiments
Our evaluation experiment was designed to answer the following four questions:
  1. 1.
    RQ1 (§ 6.2). How does Malsight performance compare across two training phases and mainstream code summarization models?
  2. 2.
    RQ2 (§ 6.3). How do module combinations and different stripping scenarios affect Malsight’s performance, particularly the annotator module?
  3. 3.
    RQ3 (§ 6.4). How does the new BLEURT-sum evaluation method compare to existing code summarization evaluation metrics?
  4. 4.
    RQ4 (§ 6.5). How does Malsight perform when applied to real-world malicious software, and how do assessments from human reverse engineers validate its usability?
6.1 Experimental Setup
Our experiment is running on the Ubuntu 20.04 system, equipped with one Intel Xeon Silver 4210 CPU 2.20 GHz, and two NVIDIA A40 GPUs with 125 GB RAM. Binary file processing tools include Radare2, IDA Pro v7.5, GCC v9.4.0, and GNU Make v4.2.1. Our programming language is Python v3.8.13, with transformers v4.16.2, torch v2.1.2+cu121.
6.2 RQ1. Performance Test
6.2.1 Code Summary Performance
During training, the model undergoes two phases. The first phase involves fine-tuning malware source code (MalS) to learn malicious behavioral features. The second phase focuses on generating high-quality code summaries (BenignC) by incorporating semantic information from pseudocode and annotations. In both phases, we split the dataset into training, cross-validation, and test sets with a ratio of 7:2:1. We then evaluated the model at each stage to ensure its effectiveness. The result is shown in Table II.
Since each phase uses a different dataset, their effectiveness should be verified independently rather than by direct score comparison. Considering there is no universally accepted threshold for model evaluation scores, we provide the performance of another state-of-the-art binary code summarization model (CP-BCS [55]) as a baseline. CP-BCS was tested on its own dataset, which consists of binaries compiled with GCC 7.3.0 for x86 architecture (32-bit) and then stripped. Our results in both phases demonstrate higher usability compared to CP-BCS.
TABLE II: Score During Two Train Phase
BLEURT-sum BLEU ROUGE-L METEOR
Phase1 74.17 18.06 38.56 24.69
Phase2 72.74 22.61 41.19 25.37
CP-BCS(Baseline) 17.78 21.50 16.89 11.92
After completing the training step, we use MalP as the test set (it is the only pseudocode summary dataset built for malware that we know until now). We compare Malsight to existing pseudocode summarization methods and the popular prompt-based general-purpose large model on dataset MalP. In the actual experiment, we built three versions of data for MalP: not-stripped, demi-stripped (only stripping the function name), and all-stripped (stripping the function name and all identifiers inside the function).
The closest version of the dataset to the real world, that is, the results of the all-stripped version as test set, are shown in Table III.
TABLE III: Comparison with baseline work
BLEURT-sum BLEU ROUGE-L METEOR AVG time (function) AVG summary length BLEURT-sum variance
BinT5 21.18 1.92 9.45 3.51 0.16 6.81 201.11
HexT5 27.67 2.71 11.23 3.22 0.13 7.11 216.81
WizardCoder-15B 53.43 7.75 23.16 13.71 2.44 32.20 303.03
Code Llama-7b 56.00 8.52 24.55 14.95 2.24 31.32 290.16
Code T5+ 17.18 1.74 4.17 2.54 0.1743 7.2869 161.34
WizardLM-2-7B 55.81 5.33 18.46 15.61 22.07 61.74 265.63
deepseek-llm-7b-chat 50.88 7.07 20.08 12.70 7.41 25.48 306.78
ChatGPT-3.5 60.09 9.96 25.19 16.54 - 25.27 296.81
Malsight (Ours) 62.14 9.80 25.11 16.87 2.51 35.27 131.65
* AVG time (function) is measured in seconds.
* AVG summary length is measured in words.
As shown in Table III, we conducted experiments using evaluation methods such as BLEURT-sum, BLEU, ROUGE, and METEOR (the usability of BLEURT-sum is substantiated in the referenced paper) to carry out experiments on Malsight, WizardLM [57], and Code Llama [58], etc. We also measured model performance using criteria such as summary length, variance, and processing time, as these factors can directly impact algorithm-based evaluations (see Appendix I for details).
As previously mentioned, general-purpose large models, trained on extensive corpora, possess broader knowledge, providing an advantage in code summarization tasks. Among these, GPT, as a commercial large model, performed the best with a BLEURT-sum score of 0.6009. Malsight’s innovative fine-tuning approach offers a significant advantage in summarizing malware code, with its annotation generation effectively bridging the knowledge gap between specialized and general-purpose models. Malsight achieved the highest scores among all methods in both BLEURT-sum and METEOR, which we consider two of the most reliable evaluation metrics.
This conclusion can be corroborated by Figure 6. The figure shows that Malsight has better score evaluation results and less variance, i.e., more stable code summary output, on test results stripped of data.
Figure 6: Data distribution. Malsight presents a relatively stable high segmentation distribution, which beats GPT-3.5 work.
6.2.2 Annotation Extraction Effect
The annotation extraction model was trained and tested on the AnnoS dataset, and the test results are shown in the table The annotated extraction model was trained and tested on the AnnoS dataset to complete the sequence SL task, dividing the data into normal codes (represented by N-label), important function APIs (represented by A-label), and important strings (represented by S-label). The model test produces the confusion matrix shown in Table IV below.
TABLE IV: Annotation extraction confusion matrix
Actual/Predicted N-label(Predicted) A-label(Predicted) I-label(Predicted)
N-label(Actual) 1,847,647 9,345 11,721
A-label(Actual) 15,865 116,099 1,784
I-label(Actual) 21,639 1,688 38,933
The results show that the accuracy of the model is 96.99%, which is almost comparable to the results obtained by manual annotation.
6.3 RQ2. Ablation Experiment
To verify the usability of each module in Malsight, we conducted ablation experiments on the annotation module and the two phases of code summarization. We continued to use MalP as a test dataset, experimenting with different module combinations.
6.3.1 Module Ablation
As shown in Table V, the modules in Malsight complemented each other effectively. This ablation study was conducted on the not-stripped version of MalP to minimize result fluctuations due to the absence of dynamic annotation.
TABLE V: Ablation experiment
Model BLEURT-sum BLEU ROUGE-L METEOR
Full Model 66.29 10.95 26.97 18.80
w/o Annotation 60.05 7.65 22.76 14.93
w/o Phase 1 62.05 10.18 24.16 17.87
w/o Phase 1 and Annotation 56.48 7.35 21.99 13.99
w/o Phase 2 61.93 9.19 26.25 15.37
w/o Phase 2 and Annotation 59.13 7.42 26.72 13.32
TABLE VI: Performance on different levels of stripping
Stripping Levels BLEURT-sum BLEU ROUGE-L METEOR AVG time (function) AVG summary length BLEURT-sum variance
Not-Stripped 66.29 10.95 26.97 18.80 2.28 32.85 111.39
Not-Stripped w/o Annotation 60.05 7.65 22.76 14.93 2.01 29.12 143.62
Demi-Stripped 64.01 10.42 26.03 17.62 2.44 34.29 123.30
Demi-Stripped w/o Annotation 58.04 8.27 23.36 14.25 2.12 30.92 146.90
All-Stripped 62.14 9.80 25.11 16.88 2.51 35.27 131.65
All-Stripped w/o Annotation 54.98 7.82 22.31 13.36 2.16 32.40 150.75
* AVG time (function) is measured in seconds.
* AVG summary length is measured in words.
We designed five experiments to test different combinations for the code summarization task, as follows: (1) Removing the annotation module and summarizing without any annotation, labeled as “w/o Annotation”; (2) Canceling the phase 1, labeled as “w/o Phase ”; (3) Canceling the phase 1 and removing the annotation module, labeled as “w/o Phase 1 and Annotatio”; (4) Canceling the phase 2, labeled as “w/o Phase 2”; (5) Canceling the phase 2 and removing the annotation module, labeled as “w/o Phase 2 and Annotation”.
The results indicate that different degrees of ablation have varying negative impacts on systematic performance. “w/o Phase 1 and Annotation” suffered a data shift and thus exhibited worse performance on the malware dataset than the other combinations, 56.48 in BLEURT-sum. Notably, the absence of the annotation generation module (static annotation generation module) significantly affects performance, demonstrating its effectiveness.
6.3.2 Annotater vs. Stripping
The influence of different levels of strip on the annotation module continues to be explored, and the results are shown in Table VI. Experiments have shown that a higher degree of stripping creates a tolerable performance degradation in the absence of annotation, demonstrating the robustness of Malsight. After the annotation module was ablated, it produced a 9.41% performance degradation (measured by a BLEURT-sum score) in the not-stripped test and nearly 12% in the all-stripped test, which proved that the annotator effectively resisted the adverse conditions caused by stripping.
Figure 7: Annotation Ablation. The Annotation generation module adds more semantic information to sentence A, making it able to output function functionality smoothly, but after ablation, it cannot.
In the generated results, we selected relatively representative sentences, as shown in Figure 7. In the yellow box, the output from the generation module without annotations shows that the large model is highly susceptible to biases due to hallucination [59]. Based on our observations, hallucinations can cause the code summary model to make false guesses about the types of parameters and internally called functions, leading to potential user misdirection. To mitigate this issue, we incorporate annotations, which not only enhance the model’s ability to summarize behavior, structure, and application scenarios but also reduce the nonsensical outputs caused by these illusions.
6.4 RQ3. Evaluation Algorithm Test
(a) BLEU
(b) METEOR
(c) word2vec
(d) MoverScore
Figure 8: Relevant method test results. The evaluation effect of two common algorithms and ML-based evaluation methods is not ideal, and their positive samples and negative samples show a certain degree of crossover.
The EvaS dataset was constructed to evaluate our code summary evaluation method alongside other popular methods. As previously mentioned, we use both positive and negative samples to test the effectiveness of these evaluation methods by assessing their ability to distinguish between the two. The performance of existing methods, illustrated in Figure 8, shows varying degrees of crossover between positive and negative samples for BLEU, METEOR, word2vec, and MoverScore. These four methods represent the current approaches in words’ overlap measure(BLEU and METEOR) and words’ embedding measure(word2vec and MoverScore), respectively.
In Figure 8, we have established two types of decision boundaries: one being orthogonal to the X-axis (i.e., distinguishing between positive and negative samples by setting a threshold), and the other being a linear function forming a slanted line. The same process is represented by BLEURT-sum as shown in Figure 9, which shows the superiority of our method.
Figure 9: BLEURT-sum performance. Compared with other methods, BLEURT-sum has significantly better ability to distinguish between positive and negative samples.
Specifically, we calculated an F1-score for these six methods (ROUGE-L is not shown in the Figure 8) on the positive and negative sample classification task of code summary sentences to measure their ability. Our method achieved an F1-score exceeding 0.9999, significantly outperforming all other evaluation methods. Among the existing methods, METEOR performed the best with an F1-score of 0.9811, while BLEU had the lowest performance at only 0.85 as the lowest. None of the other methods achieved an F1-score above 0.95. Notably, our dataset was not meticulously curated; the negative samples were entirely random sentences with different meanings, which differ from real-world scenarios where the differences in meaning might be subtler. An F1-score below 95% may therefore indicate an unacceptable tendency to misclassify. This is because a 95% F1-score means that approximately 1 in 10 samples are incorrectly judged while upholding a high recall rate. As an evaluation method, the misjudgment can be further amplified by the model results evaluated using the evaluation scheme in the following works.
6.5 RQ4. Real World Experiments
6.5.1 Human Evaluation Experiments
So far, we have evaluated Malsight’s performance at the function level. However, this approach did not allow us to fully showcase the capabilities of the dynamic annotation module. Additionally, real-world malware predominantly manifests in executable files. Hence, we manually analyzed 10 real-world malware samples and selected 79 critical functions out of them, three experienced reverse engineers add summaries to these functions based on discussion. The complete workflow of Malsight was applied to these functions and the outputs was evaluated using both BLEURT-sum and human evaluation metrics. We further calculated the variance deviation of all calculated results as a reference to demonstrate the stability of our solution and the time required for code summarization for a single malware sample.
For the human evaluation metrics, we invited ten evaluators, including five reverse engineers who are rich in experience, three experienced, and two beginners. They were asked to focus on evaluating the usability of the code summarization results and provide a score of 0, 0.5 or 1, indicating whether the summary corresponded to the original code. It is noteworthy that due to the subjectivity of human evaluation of the summaries, different evaluators may have varying opinions on the summary of the same code. Therefore, we opted to evaluate usability only and did not solicit more detailed scores beyond 0,0.5 and 1, respectively corrrepond to usable, partially usable and unusable.
6.5.2 Performance From Different Evaluators
After three sets of scores, we obtained the human assessment scores shown in Table VII. Experienced evaluators tend to give more conservative and stable scores, while less experienced evaluators give a wider range of scores with higher variance. Further, we take the arithmetic average of the scores received by these evaluators as the final human assessment score.
TABLE VII: Evaluation of Different Categories of Evaluators
Rich in experience Experienced Beginner
Score 58.14 59.51 64.73
Variance 173.12 197.49 237.44
Table b shows the mean value of BLEURT-sum and the mean value of human evaluation on 79 labeled functions. The human evaluation score of 59.87 means that most functions have exceeded 0.5, i.e. the partially usable standard.
TABLE VIII: Real Malware Performance
BLEURT-sum Human Assessment Score AVG time (file)
Score 47.22 59.87 1.90
Variance 161.61 0.06 1.21
* AVG time (per file) is measured in hours.
Figure 10: Real-world Performance. While the ratio of human assessment score to BLEURT-sum remained positive, Malsight is able to obtain higher assessment results.
Further, we formed a data point by combining a BLEURT-sum score with a human evaluation score for the same function’s code summary, thus, we plotted a distribution of 79 data points in Figure 10. It indicates that BLEURT-sum has a positive linear correlation with human evaluation metrics, which shows the rationality of BLEURT-sum. Furthermore, the data distribution indicates that, across the two distinct evaluation metrics, the majority of the evaluation scores for Malsight’s outputs are concentrated within the middle to high range, which shows that our scheme is also usable in the real world.
7 Case Study
In section RQ4, we evaluate Malsight’s performance on 10 real-world malware. To provide more direct evidence and analyze the reasons why Malsight behaves differently on different functions, three functions are presented in this section to discuss the process that Malsight goes through behind the output that is usable, the output that is partially usable, and the output that is unusable.
Figure 11: Redacted pseudocode in Case 1.
Case 1: sub_401021.
  • Malsight Summary: Set privilege, the code snippet is difficult to ascertain the function’s purpose, but it likely retrieves or obtains professional information or status.
  • Human Summary: Escalating the privileges of the virus process.
Malsight’s summary of the function shown in Figure 11 is considered usable. In this case, Malsight summary accurately identifies “setting privileges” as the key point from a lengthy function, which aligns with the task of “privilege escalation” in the human summary.
Figure 12: Redacted pseudocode in Case 2.
Case 2: sub_401D30.
  • Malsight Summary: Convert a string to a wide character, the code defines a function that performs string operations, handles exceptional conditions, and creates new objects, potentially Get the server name, the code defines a function that opens a WinHttp connection, connects to a server, and handles error messages and string operations Get the next character from the string, the code defines a function sub_401470 that performs string operations, handles exceptional conditions, and Delete the server.
  • Human Summary: Connect to a malicious domain and make a request to retrieve data using HTTP.
Malsight’s summary of the function shown in Figure 12 is considered partially usable. Malsight summary identifies that the function uses an HTTP connection to a specific server and operates on the data, which is consistent with the human summary. However, due to the lack of some background information, it fails to recognize that the specific server is malicious.
Figure 13: Redacted pseudocode in Case 3.
Case 3: TimerFunc.
  • Malsight Summary: Get the command line, the code defines a function that handles or processes data related to a hexadecimal address, retrieves information about Redis commands and modules, Call the NtTib. ExceptionList, the code defines a function that handles or processes data related to a hexadecimal address, retrieves information about Copy the setup.exe, the code defines a function that handles or processes data related to a hexadecimal address, retrieves information about Redis commands and modules Setup the setup, the code defines a function that handles or processes data related to a hexadecimal address, retrieves information about Redis commands and modules.
  • Human Summary: Checking if the autorun.ini file exists; if not, create it and configure it to auto-run setup.exe as specified in the configuration file.
Malsight’s summary of the function shown in Figure 13 is considered unusable. The function is very long, contains numerous function calls, and includes many redundant functions that hinder analysis, resulting in a poor Malsight summary.
8 Discussion
8.1 Ethics
The construction of the dataset MalS and MalP complies with GitHub’s open-source license agreements. We only used open-source repositories with GitHub open-source licenses and downloaded these open-source codes within the rate limit set by GitHub to minimize interference with GitHub’s servers to the lowest extent possible. Additionally, We enlisted volunteers to review the MalS and MalP datasets, as well as the 10 malware samples mentioned in RQ4. With the volunteers’ consent, we adopted their review results as the final dataset.
8.2 Limitations
In this section, we delve into practical issues based on our analysis of the experimental results. Additionally, we explore aspects not covered in our work and propose potential solutions.
Real-World Malware vs. Malware Function Summaries: In our exploration of real-world malware, the BLEURT-sum evaluation score of the code summary was approximately 30% lower than the experimental score obtained by MalP in Section 6.2.
The reason could be that our code summarization model lacks the ability to summarize longer functions which occasionally appear in real-world malware. Our analysis shows that about 15% of the functions in the malware contain more than 1000 tokens (while the majority of the functions in the training set stay below 300 tokens), which is likely due to different optimization configurations during compilation. These lengthy function bodies introduce a lot of information and noise, which makes it difficult for the model to extract and summarize the critical code fragments.
Therefore, to better align with real malicious code summaries, researchers should consider doing code summarization work in units of code fragments instead of functions. However, identifying and summarizing important, human-interpretable segments in them requires large amounts of labeled data, or compositional functional insights from dynamic debugging, which poses significant challenges.
Summary for assembly language rather than pseudocode: Efforts have been made to recover information from assembly code and to generate function summaries [60]. A common perspective is that decompiled assembly code suffers varying degrees of information loss during the generation of pseudocode as Intermediate Representation (IR), depending on the disassembly algorithm used. Thus, starting directly from assembly language is considered a viable solution.
In our comparison of assembly code and pseudocode representations, we found that assembly code is challenging for models pre-trained on high-level languages to understand [32]. The structural features of assembly code are completely different from those of high-level languages. Fine-tuning a model using pseudocode can confuse the model’s understanding of function-level structural features, leading to unacceptable error output, possibly due to insufficient datasets.
We train an embedding model specifically for assembly language and build a large dataset at the assembly language level for malware. This method has the potential to outperform general-purpose large models, which typically have a poor understanding of assembly code.
9 Conclusion
In this paper, we introduced Malsight, a novel framework for binary malware summarization by exploring malicious source code and benign pseudocode. The proposed Malsight involves two datasets MalS and MalP, an LLM-based summary model, and an evaluation metrics. Experimental results on three datasets show the effectiveness of the proposed framework. Future work includes the application of the proposed framework to more downstream tasks.
Appendix A Analysis of Traditional Similarity Evaluation Metric
The criteria for evaluating sentence similarity should focus on two key aspects: similar sentences should correspond to higher scores, while dissimilar sentences should correspond to lower scores. Below, we introduce some of the shortcomings of BLEU and other traditional metrics (ROUGE, METEOR) in these aspects.
A.1 Limitations of Overlap Based Methods
Problems caused by BLEU algorithm: The formula BLEU uses to calculate sentence similarity is shown in equation 17:
$BLEU=BP*exp(\sum\limits_{n=1}^{N}w_{n}\ln(Precision_{n}))$ (17)
Where $BP$ is a brevity penalty factor, $w_{n}$ is the weight of the $n$-gram, and $Precision_{n}$ is the precision of the generated candidate sentence. Specifically, equation 18 shows the construction of $Precision_{n}$, where $candidate\&reference$ is the number of overlapping occurrences between the candidate sentence and the reference sentence.
$Precision_{n}=\frac{len(candidate\&reference)+1}{len(candidate)+1}$ (18)
In this context, the Add-One Smoothing method is used to avoid zero-count problems when the n-gram size is large. However, this introduces another issue: even if the candidate sentence and reference sentence are completely unrelated, this smoothing method still produces a certain score. This effect is particularly noticeable in the case of short sentences.
We conducted experimental tests for this scenario, testing each reference-candidate pair within the sentence length range of [1,30]. Each sentence pair had zero word overlap, and our results are shown in Figure 14.
Figure 14: Bleu Score When Zero Overlap. When two sentences have no overlap, due to the structural flaws in the BLEU algorithm, they still receive a score, indicating a bias toward shorter sentences.
Even when sentence pairs are completely mismatched, BLEU scores greater than 0.3 can occur for shorter sentences. This significant deviation from reality indicates that BLEU’s scoring is distorted for short sentences in some cases.
ROUGE & METEOR, The flaw of calculating similarity in basic units of words: ROUGE and METEOR have something in common in the construction of sentence similarity evaluation algorithms. ROUGE follows the following equation 19.
$ROUGE-L=F_{LCS}=\frac{(1+\beta^{2})R_{LCS}P_{LCS}}{R_{LCS}+\beta^{2}P_{LCS}}$ (19)
Where $R_{LCS}=\frac{LCS(C,S)}{len(S)}$,$P_{LCS}=\frac{LCS(C,S)}{len(C)}$ and $\beta$ is used to give weight to recall rates. $LCS(C,S)$ is used to calculate the length of the common substring of two strings $C$ and $S$. Subjectively, when two target sentences have a higher degree of overlap in a specific word, they are given a higher ROUGE score.
METEOR designs on the basis of rouge, following equation 20, 21 and 22.
$F_{mean}=\frac{(1+\beta^{2})PR}{R+\beta P}$ (20)
$Penalty=\gamma(\frac{chunks}{unigrams\_matched})^{\theta}$ (21)
$METEOR=F_{mean}(1-Penalty)$ (22)
Where $P=\frac{n}{len(candidate)}$,$R=\frac{n}{len(reference)}$ and $n$ are the number of words where the candidate sentence and reference sentence overlap. METEOR employs exact matching, stem matching, and WordNet-based synonym matching to address the issue of words with identical meanings not being recognized as overlaps. Consequently, METEOR outperformed ROUGE in our experiments, ranking second only to BLEURT-sum. However, algorithms that rely solely on word overlap can still misjudge due to structural similarities in sentences or phrases. For instance, in code summarization, sentences might share terms like ”function”, ”aims”, or ”code,” or convey the same idea using different wording, such as ”compare two sentences” versus ”bitwise and return true/false.”
Interestingly, we found that expanding the stopword list appropriately can enhance the performance of word overlap-based methods, especially METEOR.
A.2 Limitations of Embedding Based Methods
In recent years, using word embedding-based methods such as word2vec to determine sentence similarity has become quite popular. For instance, the pseudocode for the word2vec process is illustrated in Algorithm 1.
Algorithm 1 word2vec
1:The reference sentence $ref$ and the candidate sentence $cad$
2:the similarity $score$ of $ref$ and $cad$
3:function Main($cad,ref$)
4:     cad.vector = getVector($cad$)
5:     ref.vector = getVector($ref$)
6:     $\text{score}=\frac{\text{cad}\cdot\text{ref}}{||\text{cad}||\cdot||\text{ref}||}$
7:     return score
8:end function
9:function getVector($sentence$)
10:     words = tokenize(sentence)
11:     sentence.vector = [0,0,..]
[…]
13:         sentence.vector += word.vector
14:     end for
15:     return sentence.vector
16:end function
When calculating the similarity between two sentences, word2vec adds the vectors of each word based on the tokenized results and uses the summed vector as the sentence representation. Finally, cosine similarity is employed to determine the similarity between the sentences. A direct drawback of this approach is that it loses all word order information and heavily depends on the accuracy of the embedding algorithm.
Appendix B Reverse CFG Sorting Algorithm
Algorithm 2 REsort
1:The CFG graph $G_{M}$ for a malware binary file.
2:The reverse topsort list $L_{G}$ of the function call graph.
3:function REsort($G_{M}$)
4:     for each vertex $v$ in $G_{M}.vertices$ do
5:         if $!v.seen$ then
6:              new $G_{M}^{\prime}$
7:              $G_{M}^{\prime}.vertices\leftarrow$ call Tarjan($v$, $G$)
8:              //Tarjan() traverses all strongly connected components in $G_{M}$ and contracts vertices as $G_{M}^{\prime}.vertices$.
9:         end if
10:     end for
11:     $G_{M}^{\prime}\leftarrow$ call BuildTarGraph($G_{M}^{\prime}.vertices$, $G_{M}$)
12:     //BuildTarGraph() build a new directed acyclic graph $G_{M}^{\prime}$ using $G_{M}^{\prime}.vertices$ and $G_{M}.edges$
13:     $L_{G_{M}^{\prime}}\leftarrow$ call RetopSort($G_{M}^{\prime}$)
14:     //RetopSort() obtains the reverse topological sorting sequence of $G_{M}^{\prime}$
15:     $dist[]\leftarrow$ call Dijkstra($G_{M}$)
16:     //Dijkstra() computes the multi-source shortest paths in graph $G_{M}$ from vertices with an in-degree of 0 to each vertix.
17:     $L_{G_{M}}\leftarrow[]$
18:     for each vertex $v^{\prime}$ in $L_{G_{M}^{\prime}}$ do
19:         $mindist\leftarrow\infty$, $idx\leftarrow-1$
20:         //$idx$ determines the starting vertex for DFS in each strongly connected component in $G_{M}$
21:         for each $ver$ in $v^{\prime}.subvertex$ do
22:              if $dist[ver]<mindist$ then
23:                  $mindist\leftarrow dist[ver]$, $idx\leftarrow ver$
24:              end if
25:         end for
26:         $L_{G_{M}}$.append(DFS($idx$, $v^{\prime}$))
27:         //DFS() derives a traversal order in each strongly connected component in $G_{M}$ as part of the final reverse topological sorting.
28:     end for
29:end function
TABLE IX: Performance in different training data proportion
Proportion BLEURT-sum BLEU ROUGE-L METEOR AVG summary length BLEURT-sum variance
1:1 59.12 8.98 23.04 16.14 39.83 181.01
1:2 59.81 9.15 23.50 16.42 40.75 165.51
3:4 59.53 8.68 22.95 16.39 42.21 173.39
4:3 62.14 9.80 25.11 16.88 35.27 131.65
* Due to the limited size of MalS, the model cannot be fine-tuned in the training data proportion 2:1.
* AVG summary length is measured in words.
Algorithm 2 presents our function topological sorting algorithm, with the objective of constructing the reverse topological sorting of the function call graph $G_{M}$. Due to the presence of multi-cycles (i.e., strongly connected components) in $G_{M}$, it is not feasible to perform reverse topological sorting directly. To address this, we utilize the Tarjan algorithm to contract $G_{M}$ into a new directed acyclic graph $G_{M}^{\prime}$, where reverse topological sorting can be applied to obtain an initial function traversal list $L_{G_{M}^{\prime}}$, with each node representing a strongly connected component in $G_{M}$.
Further, to achieve topological sorting for $G_{M}$, we employ the DFS algorithm to sort each node within the strongly connected components. Since our aim is to reverse-sort the entire graph $G_{M}$, with functions closer to the start function preferred towards the end of $L_{G_{M}}$, we utilize the Dijkstra algorithm to compute the multi-source shortest paths $dist[]$ from all nodes with zero in-degree to other nodes (referred to as ’multi-source’ because the function call graph $G_{M}$ may not necessarily be connected). Subsequently, by traversing $L_{G_{M}^{\prime}}$ sequentially, for each strongly connected component, the node with the minimum $dist$ value is selected as the initial node $idx$ for DFS. Leveraging the inherent stack property of DFS, $idx$ ensures its placement at the end of $L_{G_{M}}$. Finally, $L_{G_{M}}$ represents the desired function traversal order.
Appendix C Performance in different proportion of two-phase fine-tuning training data size
We sampled a total of 140,000 pieces of data in different proportions from the two datasets (MalS and BenignC) for two phases of fine-tuning. The results of the evaluation of model performance in each training data proportion are shown in the Table IX. In the experimental part of the text, we choose the ratio of 4:3 with the best effect as the actual training set ratio of Malsight.

### v3

@@
 To this end, we propose Malsight, a novel code summarization framework that can iteratively generate descriptions of binary malware by exploring malicious source code and benign pseudocode.
-Specifically, we construct the first malware summaries, MalS and MalP, using an LLM and manually refine this dataset with human effort.
-At the training stage, we tune our proposed MalT5, a novel LLM-based code model, on the MalS dataset and a benign pseudocode dataset.
+Specifically, we construct the first malware summary dataset, MalS and MalP, using an LLM and manually refine this dataset with human effort.
+At the training stage, we tune our proposed MalT5, a novel LLM-based code model, on the MalS and benign pseudocode datasets.
 Then, at the test stage, we iteratively feed the pseudocode functions into MalT5 to obtain the summary.
-Such a procedure facilitates the understanding of pseudocode structure and captures the intricate interactions between functions, thereby benefiting the usability, accuracy, and completeness of summaries.
+Such a procedure facilitates the understanding of pseudocode structure and captures the intricate interactions between functions, thereby benefiting summaries’ usability, accuracy, and completeness.
 Additionally, we propose a novel evaluation benchmark, BLEURT-sum, to measure the quality of summaries.
 Experiments on three datasets show the effectiveness of the proposed Malsight.
-Notably, our proposed MalT5, with only 0.77B parameters, delivers comparable performance to much larger ChatGPT3.5.
+Notably, our proposed MalT5, with only 0.77B parameters, delivers comparable performance to much larger Code-Llama.
 Index Terms:
 Malware, Code Summarization, Binary Code
-1 Introduction
+I Introduction
+Figure 1: The comparison of source code (left) and its pseudocode (right).
+The pseudocode includes significantly more content and a more complex structure and strips key semantic cues such as function names.
 The AV-TEST Institute [1] recently reported that over 450,000 new malicious files and potentially unwanted applications are registered daily, showing a high demand for malware understanding.
-Binary malware summarization [2] is a reverse engineering [3] task that aims to automatically generate concise human-readable descriptions of binary executable malicious files.
+Binary malware summarization [2] is a reverse engineering [3] task that aims to generate concise human-readable descriptions of binary executable malicious files automatically.
 The summarization provides security analysts with a quick understanding of the malware’s functionality and patterns when source code is unavailable, thereby benefiting a wide range of applications such as malware cracking [4] [5] [6], malware family classification [7] [8], binary code similarity detection [9] [10] [11], and large-scale malware behavior analysis [12] [13] [14].
-Figure 1: The comparison of source code (left) and its pseudocode (right).
-The pseudocode includes significantly more content and a more complex structure, and it also strips key semantic cues such as function names.
-Existing reverse engineering tools, such as IDA [15] and Ghidra  [16], can decompile executables into higher-level C-like pseudocode, while they still lack easy-to-understand semantics information.
+Existing reverse engineering tools, such as IDA [15] and Ghidra  [16], can decompile executables into higher-level C-like pseudocode while still lacking easy-to-understand semantics information.
+While tools such as IDA Pro and Ghidra have significantly advanced binary decompilation by transforming machine code into higher-level pseudocode, they still exhibit critical limitations in handling stripped binaries and entangled logic structures.
+Specifically, these tools often fail to recover semantically meaningful variable names, function identifiers, or accurate control-flow structures when binaries are stripped of symbolic information.
+This deficiency leads to pseudocode that is not only semantically impoverished but also structurally convoluted, with deeply nested branches and opaque function calls (e.g., sub_404018 instead of initLevel), as demonstrated in Fig.
+1.
+Existing small models, although lightweight, often suffer from insufficient capability to capture the complex semantics and entangled logic in stripped malware pseudocode, leading to incomplete or inaccurate summaries.
 Consequently, a line of efforts attempts to generate human-readable summaries based on pseudocode.
 Early studies rely on manual parsing or rule-based summary generation [17] [18].
-Recent large language models (LLMs), such as BinT5 [19], HexT5 [20], CodeGen [21], and WizardCoder [22], have shown great potential to produce more informative summaries.
+Recent large language models (LLMs) [19], such as BinT5 [20], HexT5 [21], CodeGen [22], and WizardCoder [23], have shown great potential to produce more informative summaries.
 However, these data-driven approaches still face critical issues, including poor usability, inaccurate explanations, and inaccurate explanations [2].
-Figure 1 shows the underlying reasons for the above issues by comparing the source code of the function “initLevel” to the corresponding pseudocode.
+Fig.
+1 shows the underlying reasons for the above issues by comparing the source code of the function “initLevel” to the corresponding pseudocode.
 We observe that the pseudocode presents 1) significantly more content that increases from 20 lines in source to 117 lines in the pseudocode, 2) a more complex and obscure structure with multi-level nesting and entangled logic.
-The pseudocode involves 29 more calls and 29 more if statements compared to the source code at the left, 3) stripping key semantic cues such as variable names and function names.
-For example, the function “initLevel” in source code is transferred to a meaningless symbol “sub_404018”.
-To address the above challenges, we present Malsight, a novel binary malware summarization framework that can iteratively generate descriptions of executable malware by exploring malicious source code and benign pseudocode.
+The pseudocode involves 29 more calls and 29 more if statements compared to the source code at the left, and 3) stripping key semantic cues such as variable names and function names.
+For example, the function “initLevel” in the source code is transferred to symbol “sub_404018”.
+To address the above challenges, we present Malsight  [^1] We will release Malsight’s source code to contribute to the community.
+, a novel binary malware summarization framework that can iteratively generate descriptions of executable malware by exploring malicious source code and benign pseudocode.
 The proposed Malsight involves three key ingredients, including a malware dataset MalS, an LLM-based malware summarization model MalT5, and an evaluation metric BLEURT-sum.
-We describe the workflow of the proposed framework in four steps as follows.
-Constructing MalS: As an LLM-based summarization model heavily relies on high-quality annotations to learn to align with domain-specific knowledge, it necessitates high-quality malware pseudocode summaries to fine-tune the LLM.
-However, the public malware pseudocode summarization dataset is unavailable so far, and building such a benchmark is quite challenging as it requires huge human involvement for accurate annotations.
-Figure 1 illustrates three challenges of understanding malware pseudocode.
-To tackle this issue, we alternatively construct MalS, a large-scale summarization dataset using an LLM model, and malicious C language source code crawled from GitHub.
+We describe the workflow of the proposed framework in four steps.
+Constructing MalS: LLM-based summarization models rely heavily on high-quality summaries to capture domain-specific knowledge.
+However, the public malware pseudocode summarization dataset is unavailable so far, and building such a benchmark is quite challenging as it requires substantial human involvement for accurate summarization, due to the challenges outlined in Fig.
+1.
+To address this, we construct MalS, a large-scale dataset with well-constructed summaries based on malicious C-language source code from GitHub.
 The proposed MalS involves nearly 90,000 malware source functions, with 20 types of malware functions.
-We also construct a small dataset MalP for testing.
-We detail such a procedure in Section 4.3.
-Training MalT5: We use CodeT5+ [23] as the foundation model of our MalT5.
-We sequentially fine-tune the proposed MalT5 model on the MalS dataset and an existing benign pseudocode summarization dataset [19].
-The underlying intuition is that the malicious semantic knowledge from malware source code summarization and function patterns from benign pseudocode summarization, which are learned from the above two datasets, respectively, can be transferred to the generation of malware pseudocode.
-By doing so, we can properly mitigate the issue of unavailable malware pseudocode summarization datasets.
-More details are available in Section 4.4.
-Performing Generation: We use an existing tool [15] to generate pseudocode of a binary file and then generate summaries using the MalT5 model.
-We first use IDA to construct the malware call graph and then develop an algorithm to transform the graph into a function list in reverse order.
-Then we iteratively fed the first function in the list to MALT5 to generate the summary.
-More details are provided in Section 4.1 and 4.2.
-Conducting Evaluation: Previous work [24] indicated that existing metrics for generation tasks, such as Bilingual Evaluation Understudy (BLEU) [25], Metric for Evaluation of Translation with Explicit ORdering (METEOR) [26], Recall-Oriented Understudy for Gisting Evaluation-Longest Common Subsequence (ROUGE-L) [27], may not well-fit for evaluation of the binary malware summarization.
+We also constructed a small dataset, MalP, for testing.
+We detail such a procedure in Section IV-C.
+Training MalT5: We propose MalT5, a model fine-tuned on both the MalS dataset and an existing benign pseudocode summarization dataset [20].
+The underlying intuition is that the malicious semantic knowledge from malware source code summarization and function patterns from benign pseudocode summarization, which is learned from the above two datasets, respectively, can be transferred to the summarization of malware pseudocode.
+On the one hand, this also effectively addresses the challenges brought about by the lengthy content and complex and obscure structure of decompiled pseudo-code mentioned in the previous text.
+On the other hand, by doing so, we can properly mitigate the issue of unavailable malware pseudocode summarization datasets.
+More details are available in Section IV-D.
+Performing Generation: Initially, IDApro generates pseudocode and extracts the call graph from a binary file.
+We then develop an algorithm to convert the call graph into a function list, preserving the original call order.
+This function list is iteratively fed into the MalT5 model to generate summaries, following the hierarchy of callers and callees.
+At the same time, we introduce static and dynamic annotation modules to supplement the lack of pseudocode semantics caused by stripping.
+In addition, in the above process, we iteratively add such annotations to maintain the semantic links between callers and callees.
+More details are provided in Section IV-A and IV-B.
+Conducting Evaluation: Previous work [24] indicated that existing metrics for generation tasks, such as Bilingual Evaluation Understudy (BLEU) [25], Metric for Evaluation of Translation with Explicit ORdering (METEOR) [26], Recall-Oriented Understudy for Gisting Evaluation-Longest Common Subsequence (ROUGE-L) [27], may not well-fit for evaluation of code summarization task.
 We thus employ BLEURT-sum, which is more sensitive to the quality of the pseudocode summary, thereby benefiting the evaluation in practice.
-More descriptions are given in Section 5.1.
-We conduct experiments on three datasets to verify the effectiveness of the proposed Malsight framework for binary malware summarization.
-The contribution of this paper can be summarized as follows [^1] We will release our Malsight to contribute to the community..
+More descriptions are given in Section V-A.
+The contribution of this paper can be summarized as follows.
   • A binary malware summarization Framework.
-We propose Malsight, a novel framework that can iteratively generate descriptions of binary malware by exploring malicious source code and benign pseudocode.
-Our MalT5 can tackle the challenges of entangled logic and stripped semantics in pseudocode.
+We introduce Malsight, a novel framework that iteratively generates descriptions of binary malware, addressing challenges like entangled logic and stripped semantics.
   • Large-scale datasets for binary malware summarization.
 We propose MalS and MalP, two novel datasets that can be used for the LLM training and testing of an LLM of binary malware summarization.
-To the best of our knowledge, the two datasets are the first in the field, involving nearly 90,000 malicious source functions and 20 types.
-Our MalS and MalP can serve as a benchmark for various binary malware understanding tasks.
-  • An LLM-based binary malware summarization model. We propose MalT5, a novel LLM for the summarization task. The proposed MalT5 is lightweight, with only 0.7B parameters.
-  • An evaluation metric for the task: We present BLEURT-sum, a novel evaluation metric that is more sensitive to the quality of pseudocode summarization.
-  • Extensive experiments.
-We conduct extensive experiments on three datasets and provide case studies to show why the proposed framework performs best among all baselines.
-Results show that our MalT5 achieves comparable performance to ChatGPT3.5.
-2 Background
-2.1 Malware Analysis Engineering
-The field of malware analysis engineering focuses on analyzing the functionality of malware by examining its binaries, typically through static analysis methods that involve observing assembly code or pseudocode [28].
-2.1.1 Binary Decompilation
+To the best of our knowledge, the two datasets are the first in the field.
+  • An LLM-based binary malware summarization model. We propose MalT5, a lightweight (0.7B parameters) LLM tailored for binary malware summarization.
+  • An evaluation metric for the task: We present BLEURT-sum, a novel evaluation metric more sensitive to pseudocode summarization quality.
+II Related Work
+II-A Malware Analysis Engineering
+The field of malware analysis engineering focuses on analyzing malware’s functionality by examining its binaries, typically through static analysis methods that involve observing assembly code or pseudocode [28].
+II-A1 Binary Decompilation
 Decompilation [29] converts executable files into human-readable pseudocode [30], which is more concise and structured than disassembled assembly code.
@@
 R2 [31], IDA [15], Ghidra [16]) and emerging methods using LLMs [32].
-However, pseudocode lacks semantic information such as function names.
-Decompiled function names are often unreadable (e.g., sub_4061C0 in IDA Pro) [33], providing a little useful pieces of information for further analysis.
-2.1.2 Human Static Analysis
+However, pseudocode lacks semantic information, such as function names.
+Decompiled function names are often unreadable (e.g., sub_4061C0 in IDA Pro) [33], providing limited pieces of information for further analysis.
+II-A2 Human Static Analysis
 In static analysis, human experts start analyzing from the function entry point [34], inferring functionality from system Application Programming Interface (API) calls, string information, and pseudocode logic.
@@
 To assist in this process, we developed Machine Learning-based (ML-based) Malsight, which optimizes and facilitates binary malware analysis.
-2.2 NLP Technologies
+II-B NLP Technologies
 In the Malsight process, we use Bidirectional Encoder Representation from Transformers (BERT) [37] and Text-to-Text Transfer Transformer (T5) [38] architecture language models to complete specific tasks.
-For the core code summary task, we build a CodeT5+ model combined with transfer learning.
-2.2.1 BERT Family
-BERT is a large-scale transformer-based language model pre-trained on a wide corpus of text using a self-supervised learning approach.
-The design of Masked Language Modeling (MLM) and Next Sentence Prediction (NSP) training tasks makes BERT perform well in the tasks of Sequence Labeling (SL), such as Named Entity Recognition (NER).
-Building on BERT, CodeBERT [39] learns code semantics through Code-Conditioned masked language modeling (CMLM) and natural language documentation generation (NLG).
-CodeBERT has been shown to perform well on code-related tasks, and since it is derived from BERT, we have reason to believe that this model can be fine-tuned to solve the problem of SL in pseudocode.
-2.2.2 T5 Family
-T5 [38], or Text-to-Text Transfer Transformer, is a sequence-to-sequence model based on the Transformer architecture that unifies various Natural Language Processing (NLP) tasks into a single framework, including text classification, question answering, summarization, translation, and text generation.
-CodeT5 [40] is an encoder-decoder model supporting code understanding and generation, built on the T5 architecture.
-It uses Natural Language-Programming Language (NL-PL) bimodal data for pre-training with identifier tagging and masked identifier prediction tasks.
-CodeT5+ [23] introduces greater architectural flexibility and additional pre-training tasks, with instruction tuning to enhance alignment with natural language instructions.
-This results in significant performance improvements on various code-related tasks.
-Previous works like HexT5 [20] and BinT5 [19] developed datasets to train models for binary code understanding, including code summarization tasks.
-These efforts demonstrate the potential of T5-based models in binary code summarization for malware analysis.
-2.2.3 Transfer Learning
-Transfer learning involves training a model on a source task with abundant labeled data to learn general features.
-When it is difficult to obtain sufficient datasets for training, transfer learning can be used to supplement them with similar or related datasets [41].
+II-B1 BERT Family
+BERT, a transformer-based model pre-trained with Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), excels in Sequence Labeling (SL) tasks like Named Entity Recognition (NER).
+CodeBERT [39], extending BERT, learns code semantics through Code-Conditioned MLM and natural language generation, showing strong performance in code tasks.
+Given its BERT origins, we explore CodeBERT for Sequence Labeling (SL) in pseudocode.
+II-B2 T5 Family
+T5 [38] is a sequence-to-sequence Transformer model that unifies NLP tasks into a single framework.
+CodeT5 [40], built on T5, extends its capabilities to code understanding and generation using NL-PL bimodal data with identifier tagging and masked identifier prediction.
+CodeT5+ [41] improves on this with instruction tuning and additional pre-training tasks, enhancing performance on code-related tasks.
+Previous works like HexT5 [21] and BinT5 [20] further demonstrate the potential of T5-based models in binary code summarization for malware analysis.
+II-B3 Transfer Learning
+Transfer learning leverages abundant labelled data from a source task to help a model learn general features and supplement insufficient training datasets with similar or related data [42].
 In Malsight, we fine-tune the CodeT5+ [41] model to achieve transfer learning from the source code summarization task to the decompiled code summarization task.
-Besides, we use dynamic and static annotation to implement feature enhancement to compensate for the poor transfer effect caused by the highly limited similarity between the source code and the stripped decompiled code.
-2.3 Code Summary Evaluation
-In code summary model evaluation, NLP text similarity algorithms compare generated results with a reference test set, replacing costly human evaluations.
-These algorithms are categorized into word overlap and word embedding measures.
-2.3.1 Words’ Overlap Measure
+II-C Code Summary Evaluation
+In code summary model evaluation, NLP text similarity algorithms, including word overlap and word embedding measures, compare generated results with reference sets, offering a cost-effective alternative to human evaluation.
+II-C1 Words’ Overlap Measure
 Early text similarity measures like BLEU [25] and ROUGE [27] rely on word n-gram overlap between generated and reference text, with BLEU focusing on precision and ROUGE on recall.
@@
 Recent work [43] highlights limitations of words’ overlap in code summary tasks. It shows that similar structures may yield high similarity scores despite differing semantics.
-2.3.2 Words’ Embedding Measure
-The words’ embedding measure evaluates semantic similarity by analyzing the distance between sentence embeddings in a vector space, often utilizing neural network learning.
+II-C2 Words’ Embedding Measure
 Word2vec [44] is a static embedding model that represents words as points in a vector space, facilitating the proximity of semantically similar words.
@@
 It achieves this by integrating diverse lexical and semantic-level supervision signals into its pre-training process and leveraging synthetic data based on pre-trained BERT, ensuring its effectiveness and versatility in various evaluation scenarios.
-Figure 2: Workflow of Malsight. The procedure involves three steps: graph traversal, annotation generation, and model summary.
-3 Motivation and Overview
-The construction of the code summary framework mainly includes annotation generation and code summary model construction, as shown in Figure 2.
-During the evaluation phase, we tested several evaluation methods and found a reasonable way to build an evaluation model for the code summary task.
-Simultaneously, our work involves the construction of multiple datasets (for subsequent stages of training and evaluation of transfer learning-based models).
-3.1 Code Summarization Process
+Figure 2: Workflow of Malsight.
+The workflow takes binary malware as input and, after going through graph traversal, annotation generation, and model summarization, iteratively produces function-level code summaries based on the function call hierarchy.
+III Motivation and Overview
+The construction of the code summary framework mainly includes three steps, as shown in Fig.
+2.
+During the evaluation phase, we test several evaluation methods and find a reasonable way to build an evaluation model for the code summary task.
+Simultaneously, our work involves the construction of multiple datasets (for subsequent stages of training and evaluation of MalT5).
+III-A Code Summarization Process
 The code summary task is split into three steps, which are function list extraction, annotation generation, and code LLM summary.
-3.1.1 Function List Extraction
-As mentioned, existing code summary methods for binary focus only on the internal information of the function.
-We introduced the call relationship between functions and worked on the entire binary as the processing unit.
-In other words, when function func_E in Figure 2 calls func_F, it is difficult for the subsequent code summarization model to correctly summarize the functionality of func_E without any information about func_F.
-(We assume that the function name of func_F has been corrupted.) Constructing a list of reverse call sequential relational functions provides a basis for the subsequent recovery of sub-functions functionality.
-3.1.2 Annotation Generation
-Iterate through the list of functions (assuming Fun_F has been processed), and Fun_E will first be added with annotations by the static annotator and the dynamic annotator, respectively.
-Fun_E uses the static annotator to obtain static annotations based on the internal information of the function code, while the dynamic annotator adds dynamic annotations based on the generated summary of the sub-function (Func_F) in the function.
-The program, in other words, sequentially restores functions according to the Control Flow Graph (CFG) diagram from the outermost to the innermost and passes function summary results inward.
-3.1.3 Code LLM Summary
+III-A1 Function List Extraction
+Unlike existing code summarization methods focusing solely on a function’s internal information, we incorporate the call relationships between functions and treat the entire binary as a processing unit.
+As Fig.
+2 shows, it is difficult for the subsequent code summarization model to correctly summarize the functionality of func_E without any information about func_F.
+(We assume that the function name of func_F has been corrupted.) Constructing a reverse-ordered function list provides a basis for the subsequent recovery of sub-functions functionality.
+III-A2 Annotation Generation
+Malsight will iteratively process the functions in the order of the reverse function list, adding annotations to each function.
+In other words, the program sequentially restores functions according to the Function Call Graph (FCG) diagram from the outermost to the innermost and passes function summary results inward.
+III-A3 Code LLM Summary
 Fun_E(annotated) is then fed into the code summary model for final code summary generation.
-In our work, we use transfer learning to adapt the model to both the functionality of malware functions and the structural features of decompiled pseudocode.
-Based on the CodeT5+ model, we have fine-tuned the code summary task.
+In our work, we use transfer learning to adapt the model to both the malware functions’ functionality and the decompiled pseudocode’s structural features.
 The tokenizer splits the code into tokens and embeddings, incorporating a self-attention mechanism into a complete vector in the encoder.
-The decoder outputs a fine-tuned prediction based on the code summary.
+The decoder outputs a fine-tuned prediction based on the code summary and passes it back to the dynamic annotator for subsequent use.
 Figure 3: A sample of existing evaluation results. Two unrelated pairs of two sentences form Sample I, and two related pairs form Sample II, but the evaluation results contradict expectations.
-3.2 Evaluation Method
-Our research has found that existing methods can not simultaneously measure the meaning, structure, word frequency, and other features of the reference sentence and the candidate sentence, so the model’s performance may be misjudged.
-Taking Figure 3 as an example, two examples show the evaluation results of BLEU, Meteor, and ROUGE-L on two pairs of real code summaries.
+III-B Evaluation Method
+We observe that the current popular text similarity measures make it difficult to evaluate the actual quality of the code summary objectively.
+Taking Fig.
+3 as an example, two examples show the evaluation results of BLEU, Meteor, and ROUGE-L on two pairs of real code summaries.
 The figure shows that two code summaries without any semantically related results in high evaluation scores (blue-framed), while two semantically similar code abstracts receive low scores (red-framed), demonstrating the shortage of existing methods.
-In the following work, we construct an ML-based code summary evaluation method BLEURT-sum by constructing a set of positive and negative samples composed of related sentence pairs and unrelated sentence pairs.
-We evaluated the usability of the model and prior art by measuring their ability to distinguish between positive and negative samples.
-3.3 Datasets Construction
-For the two core tasks mentioned above, binary code summary and code summary model evaluation, we build corresponding datasets.
-3.3.1 Dataset For Code Summary Model
-In order to avoid the data shift problem, the training of the code summary model requires a large dataset of malware pseudocode summary.
-Unfortunately, malware datasets are typically represented as collections of compiled binary files [48], with the binary code stripped, and possibly structurally confused.
-Consequently, the creation of code summary datasets for binary malware could be unfeasible without resorting to labor-intensive manual summarization.
-In this paper, our key insight is that the code summary model requires two capabilities, understanding of Malware Functionality and adaptability to disassembled pseudocode formats (including the ability to deal with annotated code).
-Therefore, We build source-based malware datasets and pseudo-code-based benign software datasets to train the model on these two capabilities separately.
-Based on Sourcefinder [49], we were able to find malware source repositories from GitHub.
-We generate descriptive labels for the extracted malware functions using a sophisticated language model, followed by manual verification and optimization.
-Meanwhile, we use the Capybara dataset provided by BinT5 [19] (a benign software dataset) to train the model adaptability to pseudocode structure.
-3.3.2 Dataset For Evaluation Model
+In subsequent work, we propose BLEURT-sum, a BLEURT-based fine-tuning algorithm designed and evaluated to be more suitable for distinguishing available code summaries.
+III-C Datasets Construction
+For the binary code summary and code summary model evaluation tasks, we build corresponding datasets.
+III-C1 Dataset For Code Summary Model
+To prevent data shift, training a code summary model requires a large dataset of malware pseudocode summaries.
+However, malware datasets [49] are usually available only as compiled binaries, often stripped and structurally complex, making it difficult to create summary datasets without labor-intensive manual efforts.
+Our critical insight is that the code summary model requires two capabilities, understanding of Malware Functionality and adaptability to disassembled pseudocode formats (including the ability to deal with annotated code).
+Therefore, we separately build datasets from malware source code obtained from GitHub [50] and benign pseudocode from the Capybara [20] dataset.
+They correspond to MalS and BenignC, respectively, where labels are generated through a carefully designed annotation process that combines large language models with human verification.
+We built the MalP dataset by compiling and decompiling 20 randomly selected repositories from the above Github malicious code repositories, resulting in 500 functions written in pseudocode for evaluation purpose.
+III-C2 Dataset For Evaluation Model
 In the evaluation phase, the evaluation method is used to measure the similarity between the model generation results and the reference results to evaluate the quality of the model generation.
-Given two sentences (generation results and the reference results) $S_{g}$ and $S_{r}$, most evaluation methods output a score $Score$ as the evaluation result.
-Therefore, if considering the use of machine learning methods, it is necessary to construct a dataset in $\{S_{g},S_{r},Score\}({Score}\in[0,1])$ format.
-The challenge is that when a dataset of $\{S_{g},S_{r}\}$ is obtained, it is a difficult job to obtain an accurate $Score$.
+Given two sentences (generation results and the reference results) $S_{g}$ and $S_{r}$, most evaluation methods output $Score$ as the evaluation result.
+Given the tremendous potential shown by machine learning methods, it is necessary to construct a dataset in the format ${S_{g},S_{r},Score}({Score}\in[0,1])$ when considering their use.
+The challenge lies in obtaining an accurate $Score$ once a dataset of ${S_{g},S_{r}}$ is constructed.
 In our subsequent work, we propose a reasonable algorithmic flow for constructing labeled datasets $EvaS$.
-3.3.3 Dataset For Static Annotater
-In annotation generation process, the static annotater includes a core information extraction module (described in detail in Section 4.2.1).
-Due to the difficulty in accurately completing the required functions using static methods, we use a machine learning model to complete the sequence labeling task of the pseudocode.
-By constructing the dataset AnnoS, we have constructed the dataset required for model training and testing.
-TABLE I: The Proposed Datasets
+III-C3 Dataset For Static Annotater
+In the annotation generation process, the static annotater includes a core information extraction module (detail in Section IV-B1).
+Due to the difficulty in accurately completing the required functions using static methods, we use a machine learning model to complete the sequence labeling(SL) task of the pseudocode.
+By constructing the dataset CSL, we have constructed the dataset required for model training and testing.
+TABLE I: The Three Proposed Datasets
 Sets for code summary model
@@
 Dataset Size(functions) Code language Anno num(avg) Usage
-AnnoS 95,000 pseudo 3.87 Train & Test
+CSL 95,000 pseudo 3.87 Train & Test
 Sets for evaluation model
@@
 EvaS 127,510 1:1 9.6 Train & Test
-To sum up, we mainly completed the construction of three sets of datasets in different application fields, as shown in Table I.
-4 Code Summarization Workflow
-Following our breakdown of the malware code summary task in Figure 2, our implementation first extracts the reverse function list, and then sequentially generates static and dynamic annotations for the items in the function list, and finally passes them into the code summary model.
-In this process, we completed the training and designing of two Domain-Specific large model, the applying of a General-Purpose large model and the implementation of several algorithms.
-we cover the implementation of each step separately in this section.
-4.1 Function List Extraction
-As mentioned earlier in Section 3.1, in the first step of the workflow, we extract the list of reverse functions from the CFG of the malware binary.
-Since the existing methods [50] do not give a completely accurate CFG extraction flow, we implement a pluggable CFG extraction module.
-It is used to provide us with a processing scheme from binary file $B_{Mal}$ to digraph $G_{M}$ as the CFG.
-By an inverse topological traversal algorithm, it is extracted from CFG in the opposite direction of the call chain, expressed as $L_{G_{M}}=[f_{1},f_{2},...,f_{n}](n=G_{M}.vertices)$, Where $f_{i}$ represents the $i$ th function of inverse topological order of $G_{M}$.
-In this study, algorithm $REsort$ was constructed, expressed as $L_{G_{M}}=REsort(G_{M})$.
-(See Appendix B for details) By applying algorithms such as Tarjan [51], Dijkstra [52], and Depth First Search (DFS), this approach successfully addressed the obstacles caused by cyclic calls and partially connected graphs in the process of generating reverse order lists.
-Function columns $L_{G_{M}}$follow the order from outside to inside in the function call diagram to ensure that the later-called function is first in the function list and is processed first in subsequent steps.
-The following process takes the function instances from $L_{G_{M}}$ in a forward order to achieve the order of restoration from the outer layer of the CFG diagram to the inner layer, specifically, recovering from the outer API call to the main function.
-4.2 Annotation Generation
+To sum up, we construct three datasets to address the different tasks arising from the various stages of the Malsight framework, as shown in Table I.
+IV Code Summarization Workflow
+Following our breakdown of the malware code summary task in Fig. 2, we construct comprehensive systems incorporating multiple LLMs and algorithms.
+IV-A Function List Extraction
+In the first module of Malsight, we extract a reverse function list from the malware’s FCG.
+Since existing methods [51] don’t fully ensure accurate FCG extraction, we implement a pluggable module to map the binary $B_{Mal}$ to the FCG digraph $G_{M}$.
+Using an inverse topological traversal algorithm, $REsort$, we generate the reverse order list $L_{G_{M}}$, where functions are arranged according to their inverse topological order.
+Algorithm 1 REsort
+1:The FCG graph $G_{M}$ for a malware binary file.
+2:The reverse topsort list $L_{G}$ of the function call graph.
+3:function REsort($G_{M}$)
+4:     for each vertex $v$ in $G_{M}.vertices$ do
+5:         if $!v.seen$ then
+6:              new $G_{M}^{\prime}$
+7:              $G_{M}^{\prime}.vertices\leftarrow$ call Tarjan($v$, $G$)
+8:         end if
+9:     end for
+10:     $G_{M}^{\prime}\leftarrow$ call BuildTarGraph($G_{M}^{\prime}.vertices$, $G_{M}$)
+11:     $L_{G_{M}^{\prime}}\leftarrow$ call RetopSort($G_{M}^{\prime}$)
+12:     $dist[]\leftarrow$ call Dijkstra($G_{M}$)
+13:     $L_{G_{M}}\leftarrow[]$
+14:     for each vertex $v^{\prime}$ in $L_{G_{M}^{\prime}}$ do
+15:         $mindist\leftarrow\infty$, $idx\leftarrow-1$
+16:         for each $ver$ in $v^{\prime}.subvertex$ do
+17:              if $dist[ver]<mindist$ then
+18:                  $mindist\leftarrow dist[ver]$, $idx\leftarrow ver$
+19:              end if
+20:         end for
+21:         $L_{G_{M}}$.append(DFS($idx$, $v^{\prime}$))
+22:     end for
+23:end function
+Algorithm 1 shows that equipped with Tarjan, Dijkstra, and DFS, our proposed algorithm is able to address challenging issues such as cyclic calls and partially connected graphs.
+The function list $L_{G_{M}}$ is processed from outer API calls to the main function to ensure the correct restoration order.
+Specifically, $L_{G_{M}}$ follows the call hierarchy from the outermost to the innermost layers, ensuring that functions called later appear earlier in the list and are processed first in subsequent steps.
+IV-B Annotation Generation
 Figure 4: Annotation workflow.
@@
 Func’s code summary is then passed to the Caller as a dynamic annotation.
-In the order of traversal provided by the reverse function list, annotations are added to each function in turn to provide richer information. Annotations can be divided into static and dynamic types.
-We have observed that attackers frequently employ techniques like stripping to hinder reverse engineering efforts.
-This process removes crucial symbolic information, such as identifier names, from the binaries, leading to significant semantic loss in the generated pseudocode.
-As a result, code summary models for malware face substantial challenges, and the performance of code summarization tasks is adversely affected.
-To address this issue, we propose utilizing dynamic and static annotation to supplement the semantic information and enhance the features of the stripped pseudocode as Figure 4 shows.
-This approach aims to compensate for the poor transferability caused by the stark dissimilarity between the source code and the stripped decompiled code.
-4.2.1 Static Annotation
-Based on our long-term exploration of pseudocode for stripped malware, we deem that although the stripped pseudocode has a serious semantic loss, some extremely critical API calls (such as operating system APIs) and some special forms of strings that are preserved after stripping provide us with ideas for behavior analysis and semantic recovery of malware.
-Consequently, we consider building a static annotation module to provide additional information for subsequent code summaries.
-In general, the static annotation process can be divided into three parts: sequence labeling, online retrieval and annotation generation.
-Sequence labeling model: By manually labeling approximately 300,000 tokens within nearly 80,000 functions, we construct a labeled dataset $Set_{SL}$ based on $BenignS$ to train the sequence labeling model.
+In $L_{G_{M}}=\{f_{1},f_{2}...f_{len(L_{G_{M}})}\}$, we sequentially add static and dynamic annotations to enrich the information for each function $f_{i}$.
+As mentioned, attackers often use techniques like stripping to remove symbolic information, leading to significant semantic loss in the pseudocode and hindering code summarization models.
+To address this, we propose using dynamic and static annotations to supplement the missing semantics and enhance pseudocode features.
+This approach, as illustrated in Fig.
+4, helps bridge the gap between source code and stripped pseudocode.
+IV-B1 Static Annotation
+Through our extensive study of stripped malware pseudocode, we identified that critical API calls (e.g., OS APIs) and certain preserved strings offer valuable clues for malware behavior analysis and semantic recovery despite the significant semantic loss.
+Thus, we design a static annotation module to enhance code summaries, which operates in three key steps: sequence labeling, online retrieval, and annotation generation.
+Sequence labeling model: By manually labeling approximately 300,000 tokens within nearly 95,000 functions with the assistance of lex [52], we construct a labeled dataset CSL based on BenignC to train the sequence labeling model.
 Formally, the function $f_{i}$ is first sliced into an n-token code sequence by the tokenizer $T$.
@@
 Equation (1) formalizes this dataset.
-$Set_{CSL}=\{d_{0},d_{1},...,d_{N}\}$ (1)
-Subsequently, we opted to utilize the CodeBERT [39] model for training the sequence labeling task using this dataset CSL.
-$B$ is the CodeBERT base model, and $C$ is the classifier head. As a sequence is inputted, the complete model outputs the predicted labels of its tokens. Equation (2) and (3) formalize this.
-$B(s_{i})=\{o_{0},o_{1},...,o_{n-1}\}\vspace{-0.2cm}$ (2)
+$CSL=\{d_{0},d_{1},...,d_{N}\}$ (1)
+Subsequently, we utilize the CodeBERT [39] model for training the SL task using this CSL.
+As shown in equation (2) and (3), $B$ is the CodeBERT base model, and $C$ is the classifier head.
+As a sequence $s_{i}$ is inputted, the complete model outputs the predicted labels $y_{i}$ of its tokens $o_{i}$.
+$B(s_{i})=\{o_{0},o_{1},...,o_{n-1}\}\vspace{-0.1cm}$ (2)
 $C(o_{i})=y_{i}$ (3)
-We further formalize the target function in equation (4), where $l_{i_{c}}$ is the truth label and $y_{i_{c}}$ is the Softmax probability for the $c^{th}$ class.
+Further, the target function is formalized in equation (4), where $l_{i_{c}}$ is the truth label and $y_{i_{c}}$ is the Softmax probability for the $c^{th}$ class.
 $LF=-\sum\limits_{c=0}^{2}l_{i_{c}}\log{y_{i_{c}}}$ (4)
@@
 This necessitates that the model effectively classifies code-tokens to accurately label the key API calls and special strings within the pseudocode.
-Label to annotation: Once we can get the key API calls and special strings in the stripped pseudocode, we use GitHub Code Search [53] to retrieve the relevant context in the GitHub repositories.
-In the implementation, we keep the first three blocks of the search results (this is because the code search has already sorted the relevance of the results [54]).
-The outcome of random sampling and manual discrimination reveals that approximately 54.8% of the function contexts contain code comments closely associated with the functionality of the function.
-For the remaining functions, nearly 90% also offer contextual information related to the function’s operation, such as parameter names, interconnected functions, and processing logic.
-Only a small number of functions yield invalid search results.
-The filtered and preprocessed code snippets will be continuously input into the prompt-based generic model for generating static annotation.
-4.2.2 Dynamic Annotation
-As shown in Figure 4 (the blue parts represent the steps in which the annotation was added), the summary of the callee is provided to the caller as a complement to the semantic information, which we define as dynamic annotation.
-This is consistent with the actual analysis flow of binary malware analysis by reverse workers, i.e., analyzing the call relationship from the inner layer of the CFG diagram to the outer layer (corresponding function list generated in Section 4.1) and summarizing the function from the outside in (corresponding to the passing of dynamic annotation).
-In this way, we can make full use of the dynamic behavior characteristics implied by the call relationships between functions in the pseudocode.
-4.3 Building Malware Datasets
+Label to annotation:
+Once we obtain the key API calls and special string literals from the stripped pseudocode, we utilize GitHub Code Search [53] to retrieve relevant contextual information from public GitHub repositories.
+In practice, we construct a query based on the extracted identifiers and submit it to the code search engine.
+To ensure relevance while maintaining efficiency, we retain only the top three retrieved blocks for each query.
+This decision is motivated by the observation that GitHub Code Search sorts results by estimated relevance according to their ranking algorithm [54], and that the top-ranked blocks typically contain the most pertinent information.
+We conducted random sampling and manual inspection across the retrieved results.
+The analysis shows that approximately 54.8% of the retrieved function contexts contain comments explicitly describing the function’s intended functionality.
+Among the remaining functions, nearly 90% still provide valuable contextual signals, such as meaningful parameter names, semantically related helper functions, control flow hints, and call dependencies.
+Only a small minority of cases yield irrelevant or unusable retrievals, often due to noise in the search index or ambiguous identifiers.
+After retrieval, the code snippets undergo a filtering and preprocessing step.
+Specifically, we discard results that are excessively short, semantically incomplete, or duplicate the input pseudocode without adding new context.
+The remaining high-quality snippets are then incrementally injected into the prompt of a generic, instruction-following large language model to generate static natural language annotations for each function, denoted by $AnnoS(f_{i})$.
+This retrieval-augmented annotation strategy significantly enriches the available context for static analysis and serves as an important intermediate process in our overall framework.
+IV-B2 Dynamic Annotation
+As shown in Fig.
+4 (the blue parts represent the steps in which the annotation was added), the summary of the callee is provided to the caller as a complement to the semantic information, which we define as dynamic annotation.
+This is consistent with the actual analysis flow of binary malware analysis by reverse workers, i.e., analyzing the call relationship from the inner layer of the FCG diagram to the outer layer (corresponding function list generated in Section IV-A) and summarizing the function from the outside in (corresponding to the passing of dynamic annotation).
+In this way, we can make full use of the dynamic behavior characteristics implied by the call relationships between functions in the pseudocode, denoted by $AnnoD(f_{i})$.
+The function $f_{i}=f_{i}\oplus AnnoS(f_{i})\oplus AnnoD(f_{i})$ will be passed into the code summary model for later processing.
+IV-C Building Malware Datasets
 In the traditional scheme of building datasets for decompiled code, the datasets are built at the function level [55].
@@
 $Set_{ideal}=\{(f_{n}^{pse},f_{n}^{sum})|f_{n}^{sum}=SUM(f_{n}^{pse})\}$ (5)
-Out of 2,289 GitHub repositories that were determined to be malware, we extracted close to 30K functions.
-We filter for functions that repeat, shorter than five lines, and format-challenged functions, resulting in a dataset of 89,609 functions.
+Out of 2,289 GitHub repositories that are determined to be malware, we extract close to 30K functions.
+We filter out functions that are repetitive, shorter than five lines, or have formatting issues, resulting in a dataset of 89,609 functions.
 (The lack of strict filtering may lead to overlap between the train and test sets, consequently yielding inflated results.)
-Unlike benign open-source projects that are well maintained, most of the functions we extract do not have comments in context for us to label as code summaries.
-Fortunately, the semantic information in the source code is rich enough that we used a well-designed prompt to complete the code summary for us via GPT3.5-Turbo [56].
-In this way, we extract dataset MalS in equation (6) format.
-Further, we build a 500-function dataset MalP for testing our model.
-The reason why dataset MalP has a relatively small quantity is that it was obtained by manually compiling and decompiling the git repository.
-MalP is compiled from the makefile provided by the developer, so we have not mentioned the configuration related to compilation optimization.
+Unlike well-maintained open-source projects, most extracted functions lack comments for use as code summaries.
+However, the source code’s rich semantics allowed us to generate summaries using a well-designed prompt with LLM, and engaged five human experts to assist in reviewing and refining the generated summaries, resulting in the MalS dataset in equation (6).
+We built the MalP dataset with 500 functions, selected from 20 randomly chosen repositories.
+Its smaller size is due to the difficulty of manually compiling and decompiling poorly maintained malware repositories, from which we picked functions with strong functionality characteristics.
 $Set_{MalS}=\{(f_{n}^{Sou},f_{n}^{sum})|f_{n}^{sum}=SUM(f_{n}^{Sou})\}$ (6)
 $Set_{MalP}=\{(f_{n}^{pse},f_{n}^{sum})|f_{n}^{sum}=SUM(f_{n}^{Sou})\}$ (7)
-As mentioned above, the malware training set we built was made up of source code, so another dataset was needed to help our model understand the stripped function features.
-In the transfer learning option, we used the Capybara dataset (a benign software dataset) provided by BinT5 [19] and annotated it to provide our model with adaptations to the annotated code summary task.
-It is provided as $Set_{Capybara}$, and we process it as $Set_{benignC}$, where the $ANN()$ is the static annotation generation method to form PL-NL bimodal data $f_{n}^{ann}$.
+Since MalS is a source code training set, we need an additional dataset to help our model understand stripped pseudocode features.
+The Capybara dataset provided by BinT5 [20] is a benign pseudocode dataset.
+We process the Capybara dataset ($Set_{Capybara}$) using our system’s static annotation module ($ANNS()$), transforming it into BenignC.
 $Set_{Capybara}=\{(f_{n}^{pse},f_{n}^{sum})|f_{n}^{sum}=SUM(f_{n}^{Sou})\}$ (8)
-$Set_{benignC}=\{(f_{n}^{ann},f_{n}^{sum})|f_{n}^{ann}=ANN(f_{n}^{pse})\}$ (9)
-In the process of training the code summary model, we use $Set_{MalS}$ and $Set_{benignC}$ to complete the transfer learning process, and $Set_{MalP}$ to test during the evaluation phase.
-Figure 5: Code Summary Topic Distribution. The results of topic analysis for ${Set_{MalS}}$ indicate that its distribution conforms to the functional distribution of malware functions.
-For $Set_{MalS}$, 25% of the samples were extracted for thematic analysis of function functionality, as shown in Figure 5.
-Our analysis results confirm that $Set_{MalS}$ and $Set_{BenignC}$ have significantly different theme distributions.
+$Set_{benignC}=\{(f_{n}^{ann},f_{n}^{sum})|f_{n}^{ann}=ANNS(f_{n}^{pse})\}$ (9)
+When training the code summary model, we use MalS and BenignC to complete the transfer learning tuning process, and $Set_{MalP}$ to test during the evaluation phase.
+Figure 5: Topic distribution of code summary. The results of topic analysis for ${Set_{MalS}}$ indicate that its distribution conforms to the functional distribution of malware functions.
+Additionally, 25% samples from MalS were randomly extracted for thematic analysis of function functionality, as shown in Fig.
+5.
+Our analysis results confirm that MalS and BenignC have significantly different theme distributions.
 Among a large number of security-related functions in the former, there are an unignorable number of codes for lock, permissions, and encryption, while in the latter, there are a large number of codes related to game logic and driver calls which irrelevant to malware.
-4.4 Code Summary Model
+IV-D Code Summary Model
 In Malsight, we fine-tune the CodeT5+ model to enable transfer learning from source code summarization to decompiled pseudocode summarization.
-To achieve this, we have designed two distinct phases of fine-tuning to facilitate smooth transfers that accommodate variations in data characteristics and distribution biases across different datasets.
-Phase 1.
-In the initial phase, we fine-tune the model that uses the dataset we built consisting of malware source code $Set_{MalS}$.
-The source code of malware has a similar semantic structure to the code in ordinary scenarios, but it also has behavioral or semantic features that the latter does not exist, which are represented by some code fragments with malicious purposes, such as self-replication and propagation, illegal access to system resources, and vulnerability exploitation.
-Through this phase of fine-tuning, the model can learn the behavioral or semantic features of these malicious codes.
-In this phase, the poor effects of data shift are addressed by using $Set_{MalS}$ with the same functional distribution as malware in the real world.
-Phase 2.
-In this phase, we ask the model to learn the corresponding semantic information from the pseudo-code and the annotation text simultaneously to better assist the model in generating high-quality code summaries.
-We use the PL-NL bimodal dataset BenignC to fine-tune the encoder of the model which we fine-tuned in the previous phase in case the decoder’s parameters are frozen.
-By only fine-tuning the encoder, we allow the model to adjust itself to better suit the specific task at hand, without updating the weights too much.
-This can not only lead to a more robust and generalized understanding of the data but also shorten the training time due to the reduction in parameters.
-Especially, the raw input is processed into a standardized format as equation (10) where $t_{c_{i}}$ is the tokens from a code sequence, $t_{a_{i}}$ is the tokens from the corresponding annotation, $t_{sep}$ is a special token in CodeT5+ to separate the inputs in different modes.
-Inputs in such a format can assist the encoder in distinguishing the difference between two modes and learning the correlation between them.
-$f_{n}^{ann}\to\{t_{c_{0}},...,t_{c_{n-1}},t_{sep},t_{a_{0}},...,t_{a_{m-1}}\}$ (10)
-After the above two phases of fine-tuning, the model basically has the ability to accept bimodal input composed of pseudocode and annotation text and generate code summary.
-We use this model to summarize the entire malware in reverse of the function call order.
-5 Evaluation Method
-As mentioned in Section 3.3.2, the evaluation algorithm for code summary tasks should accept the reference as input and separate available and unavailable generated results.
-In this section, we introduce our exploration of code summary dataset construction (EvaS) and evaluation model construction (BLEURT-sum) respectively.
-5.1 Evaluation Dataset Construction
-Utilizing the code summary from the MalS dataset as a foundation for our research, we curated a positive and negative sample pair for the tuning of our new evaluation model BLEURT-sum.
-The positive sample consisted of two code summary result sentences sharing the same meaning and was initialized as $\{S_{g},S_{r},1\}$, while the negative sample comprised two randomly different sentences and was initialized as $\{S_{g},S_{r},0\}$.
-$S_{g}$ and $S_{r}$ represent generated statement and reference statement, respectively.
-In order to build the dataset in $\{S_{g},S_{r},Score\}({Score}\in[0,1])$ format, one possible idea is to build an algorithm, $Score=GenSim(S_{g},S_{r},0or1)$, to automatically generate the label $Score$ required for model training.
-Therefore, the key task is to construct a reliable $GenSim()$ function.
-When $S_{g}==SUM(f_{n})$, contrusted $Score$ can be expressed using equation (11), considering that the SUM() function produces different outputs when faced with the same input, which is a non-deterministic function.
+To achieve this, we design two distinct phases of fine-tuning to facilitate smooth transfers that accommodate variations in data characteristics and distribution biases across different datasets.
+In Malsight, we leverage transfer learning to shift the model from source code summarization to decompiled pseudocode summarization.
+This is achieved by utilizing two datasets: MalS (source code dataset) and BenignC (pseudocode dataset), aiming to approximate the ideal training effect on an unachievable dataset $Set_{ideal}$.
+First, the model is trained on MalS for semantic features of the malware source code.
+For each sample $(f_{n}^{Sou},f_{n}^{sum})$ in MalS, where $f_{n}^{Sou}$ represents a source code snippet and $f_{n}^{sum}$ represents the corresponding summary, the model minimizes the loss function as equation (10).
+Here, $M_{\theta}$ denotes the model parameters, and $\mathcal{L}$ represents the loss function (e.g., cross-entropy loss).
+The output of this stage is the encoded feature representation of the source code $h_{MalS}=\text{Encoder}_{\theta}(x_{i})$.
+$L_{MalS}=\frac{1}{n_{MalS}}\sum_{i=1}^{n_{MalS}}\mathcal{L}(M_{\theta}(f_{n}^{Sou}),f_{n}^{sum})$ (10)
+Next, the model undergoes transfer learning using BenignC (pseudocode dataset) to further refine its capabilities.
+For each sample $(f_{i}^{ann},f_{n}^{sum})$ in BenignC, where $f_{i}^{ann}$ is the pseudocode sequence and $f_{n}^{sum}$ is the annotation, the data processing pipeline is shown in equation (11).
+$f_{i}^{ann}\to\{fc_{0}^{ann},\dots,fc_{p-1}^{ann},f_{sep},fc_{0}^{ann},\dots,fa_{q-1}\}$ (11)
+The input pseudocode and annotation are concatenated into a bimodal sequence where $fc_{p}$ represents tokens from the pseudocode, $fa_{q}$ represents tokens from the annotation, and $f_{sep}$ is a special separator token.
+The T5 model’s encoder processes this sequence of pseudocode and annotations, generating an encoded representation $h_{benignC}=\text{Encoder}_{\theta}(f_{i}^{ann})$.
+This representation serves as the basis for generating the corresponding code summary.
+To optimize the model’s performance, we minimize the loss function defined in equation (12), where $M_{\theta}$ represents the parameters of the model, $f_{i}^{ann}$ is the input pseudocode sequence, and $f_{i}^{sum}$ is the target summary.
+$L_{benignC}=\frac{1}{n_{benignC}}\sum_{i=1}^{n_{benignC}}\mathcal{L}(M_{\theta}(f_{i}^{ann}),f_{i}^{sum})$ (12)
+This loss function captures the difference between the predicted and target summaries across the entire dataset.
+The model leverages transfer learning by training on both MalS (source code) and BenignC (pseudocode), allowing it to learn generalized features from source code and apply them to pseudocode summarization.
+By facilitating knowledge transfer between these datasets, the T5 encoder ensures consistency and improves its ability to handle different data modalities, enhancing performance on both source code and pseudocode tasks.
+$L_{total}=\alpha L_{MalS}+\beta L_{benignC}$ (13)
+The overall objective loss for the transfer learning process is defined in equation (13).
+Specifically, $\alpha$ and $\beta$ are balancing factors between the two learning objectives.
+This method allows the model to adapt to the semantic structure of pseudocode and generate high-quality code summaries.
+V Evaluation Method
+As mentioned in Section III-C2, the evaluation algorithm for code summary tasks takes the reference as input and distinguish between valid and invalid generated results.
+This section introduces our exploration of EvaS and BLEURT-sum, respectively.
+V-A Evaluation Dataset Construction
+Utilizing the code summary from the MalS dataset as a foundation for our research, we curate a positive and negative sample pair for the tuning of our new evaluation model BLEURT-sum.
+The positive sample consisted of two code summary result sentences sharing the same meaning is initialized as $\{S_{g},S_{r},1\}$, while the negative sample comprised two randomly different sentences and is initialized as $\{S_{g},S_{r},0\}$.
+$S_{g}$ and $S_{r}$ represent generated statements and reference statements, respectively.
+To build the dataset in $\{S_{g},S_{r},Score\}({Score}\in[0,1])$ format, one approach is to build an algorithm, $Score=GenSim(S_{g},S_{r},0or1)$, which automatically generate the $Score$ labels needed for model training.
+The key task is constructing a reliable $GenSim()$ function.
+When $S_{g}==SUM(f_{n})$, the $Score$ can be expressed using equation (14).
+This accounts for the fact that the SUM() function may produce different outputs for the same input, as it is a non-deterministic function.
 $S_{r}\!=\!\text{SUM}(f_{n})\!\implies\!Score\!=\!\begin{cases}1\!&\!\text{if }S_{g}=\text{SUM}(f_{n})\\ 0\!&\!\text{if }S_{g}=\text{SUM}(\neg f_{n})\end{cases}$ (14)
-Since when $\{S_{g},S_{r}\}$ pairs were built, no sentence structure dependencies are taken into account, which means the equation (11) of Score can contain a few words overlap-based features (§ 2.3), the natural consideration is to combine the characteristics of sentence structure and semantics.
-Our idea is to solve the proportion of semantic information and sentence structure information in sentence similarity evaluation.
-Taking the original 0,1 tag as the semantic feature, we further extract the static feature to get a Score, multiply the two by the corresponding proportion respectively, and then add them to get $Score$.
+When constructing the $\{S_{g},S_{r}\}$ pairs, sentence structure dependencies are not considered, meaning that the score calculated in equation (14) is based purely on semantic information.
+To enhance this, we propose incorporating both sentence structure and semantics into the evaluation process.
+Our approach seeks to determine the appropriate balance between these two aspects in sentence similarity evaluation.
+Taking the original 0,1 tag as the semantic feature, we further extract the static feature to get $s_{f}$.
 Assuming that the proportion of semantic information is $p$, the structural feature calculation function is $Struc()$, the following equations formalize this Score construction method.
-$s_{f}=Struc(S_{g},S_{g})$ (12)
+$s_{f}=Struc(S_{g},S_{r})$ (15)
 $Score\!=\!\begin{cases}p+(1-p)*s_{f}\!&\!\text{if}S_{g}=\!\text{SUM}(f_{n})\\ (1-p)*s_{f}\!&\!\text{if}S_{g}=\!\text{SUM}(f_{\neg n})\end{cases}$ (16)
@@
 Then we solve the minimum square error and get the semantic and structural ratio close to 1:4, which makes $p=1/5$.
-5.2 Evaluation Model
-Based on our insights into the evaluation model, it should be able to combine both structural and semantic features and give an evaluation score for the specific task of code summary.
+V-B Evaluation Model
+Based on our insights of the evaluation model, it should be able to combine both structural and semantic features and give an evaluation score for the specific task of code summary.
 BLEURT is based on BERT and adds additional pre-training steps between pre-training and fine-tuning to the synthesized data.
@@
 The rich training signals ensure the universality of BLEURT and the comprehensiveness of the evaluation angle.
-We then further fine-tune the BLEURT model on the code summary sentence pair dataset.
+We then further fine-tune the BLEURT model on EvaS dataset.
 First, the generated text and the reference text are input together into the model and the vector is generated as shown in equation (17).
-This solution is called BLEURT-sum.
 $v_{[CLS]},v_{S_{g1}},...,v_{S_{gn}},...,v_{S_{rn}}=BLEURT(S_{g},S_{r})$ (17)
@@
 $loss=\frac{1}{N}\sum\limits_{n=1}^{N}\left|\left|Score-\hat{Score}\right|\right|^{2}$ (19)
-In existing works, Word overlap measures are commonly used in text similarity evaluation but perform poorly for code summary evaluation.
-Code summaries often include keywords like “retur” and “initializ”, which do not significantly contribute to the overall meaning.
-These keywords can inflate overlap rates and lead to misleadingly high scores.
-Therefore, BLEURT-sum on the one hand still considers Word overlap measures, and on the other hand introduces more dimensions, which greatly improves the performance in code summary evaluation tasks.
-6 Experiments
-Our evaluation experiment was designed to answer the following four questions:
-  1. 1.
-    RQ1 (§ 6.2). How does Malsight performance compare across two training phases and mainstream code summarization models?
-  2. 2.
-    RQ2 (§ 6.3). How do module combinations and different stripping scenarios affect Malsight’s performance, particularly the annotator module?
-  3. 3.
-    RQ3 (§ 6.4). How does the new BLEURT-sum evaluation method compare to existing code summarization evaluation metrics?
-  4. 4.
-    RQ4 (§ 6.5). How does Malsight perform when applied to real-world malicious software, and how do assessments from human reverse engineers validate its usability?
-6.1 Experimental Setup
-Our experiment is running on the Ubuntu 20.04 system, equipped with one Intel Xeon Silver 4210 CPU 2.20 GHz, and two NVIDIA A40 GPUs with 125 GB RAM.
-Binary file processing tools include Radare2, IDA Pro v7.5, GCC v9.4.0, and GNU Make v4.2.1.
-Our programming language is Python v3.8.13, with transformers v4.16.2, torch v2.1.2+cu121.
-6.2 RQ1. Performance Test
-6.2.1 Code Summary Performance
-During training, the model undergoes two phases.
-The first phase involves fine-tuning malware source code (MalS) to learn malicious behavioral features.
-The second phase focuses on generating high-quality code summaries (BenignC) by incorporating semantic information from pseudocode and annotations.
-In both phases, we split the dataset into training, cross-validation, and test sets with a ratio of 7:2:1.
-We then evaluated the model at each stage to ensure its effectiveness.
-The result is shown in Table II.
-Since each phase uses a different dataset, their effectiveness should be verified independently rather than by direct score comparison.
-Considering there is no universally accepted threshold for model evaluation scores, we provide the performance of another state-of-the-art binary code summarization model (CP-BCS [55]) as a baseline.
-CP-BCS was tested on its own dataset, which consists of binaries compiled with GCC 7.3.0 for x86 architecture (32-bit) and then stripped.
-Our results in both phases demonstrate higher usability compared to CP-BCS.
-TABLE II: Score During Two Train Phase
-BLEURT-sum BLEU ROUGE-L METEOR
-Phase1 74.17 18.06 38.56 24.69
-Phase2 72.74 22.61 41.19 25.37
-CP-BCS(Baseline) 17.78 21.50 16.89 11.92
-After completing the training step, we use MalP as the test set (it is the only pseudocode summary dataset built for malware that we know until now).
-We compare Malsight to existing pseudocode summarization methods and the popular prompt-based general-purpose large model on dataset MalP.
-In the actual experiment, we built three versions of data for MalP: not-stripped, demi-stripped (only stripping the function name), and all-stripped (stripping the function name and all identifiers inside the function).
-The closest version of the dataset to the real world, that is, the results of the all-stripped version as test set, are shown in Table III.
-TABLE III: Comparison with baseline work
+In summary, existing word overlap methods often inflate scores due to common keywords like ”return” or ”initialize”, which may be incremental.
+Meanwhile, the BLEURT-sum improves performance by considering both word overlap and additional dimensions.
+VI Experiments
+TABLE II: Comparisons with baseline work
 BLEURT-sum BLEU ROUGE-L METEOR AVG time (function) AVG summary length BLEURT-sum variance
@@
 deepseek-llm-7b-chat 50.88 7.07 20.08 12.70 7.41 25.48 306.78
-ChatGPT-3.5 60.09 9.96 25.19 16.54 - 25.27 296.81
 Malsight (Ours) 62.14 9.80 25.11 16.87 2.51 35.27 131.65
@@
 * AVG summary length is measured in words.
-As shown in Table III, we conducted experiments using evaluation methods such as BLEURT-sum, BLEU, ROUGE, and METEOR (the usability of BLEURT-sum is substantiated in the referenced paper) to carry out experiments on Malsight, WizardLM [57], and Code Llama [58], etc.
-We also measured model performance using criteria such as summary length, variance, and processing time, as these factors can directly impact algorithm-based evaluations (see Appendix I for details).
-As previously mentioned, general-purpose large models, trained on extensive corpora, possess broader knowledge, providing an advantage in code summarization tasks.
-Among these, GPT, as a commercial large model, performed the best with a BLEURT-sum score of 0.6009.
-Malsight’s innovative fine-tuning approach offers a significant advantage in summarizing malware code, with its annotation generation effectively bridging the knowledge gap between specialized and general-purpose models.
-Malsight achieved the highest scores among all methods in both BLEURT-sum and METEOR, which we consider two of the most reliable evaluation metrics.
-This conclusion can be corroborated by Figure 6.
-The figure shows that Malsight has better score evaluation results and less variance, i.e., more stable code summary output, on test results stripped of data.
-Figure 6: Data distribution. Malsight presents a relatively stable high segmentation distribution, which beats GPT-3.5 work.
-6.2.2 Annotation Extraction Effect
-The annotation extraction model was trained and tested on the AnnoS dataset, and the test results are shown in the table The annotated extraction model was trained and tested on the AnnoS dataset to complete the sequence SL task, dividing the data into normal codes (represented by N-label), important function APIs (represented by A-label), and important strings (represented by S-label).
-The model test produces the confusion matrix shown in Table IV below.
-TABLE IV: Annotation extraction confusion matrix
-Actual/Predicted N-label(Predicted) A-label(Predicted) I-label(Predicted)
-N-label(Actual) 1,847,647 9,345 11,721
-A-label(Actual) 15,865 116,099 1,784
-I-label(Actual) 21,639 1,688 38,933
-The results show that the accuracy of the model is 96.99%, which is almost comparable to the results obtained by manual annotation.
-6.3 RQ2. Ablation Experiment
-To verify the usability of each module in Malsight, we conducted ablation experiments on the annotation module and the two phases of code summarization.
-We continued to use MalP as a test dataset, experimenting with different module combinations.
-6.3.1 Module Ablation
-As shown in Table V, the modules in Malsight complemented each other effectively.
-This ablation study was conducted on the not-stripped version of MalP to minimize result fluctuations due to the absence of dynamic annotation.
-TABLE V: Ablation experiment
-Model BLEURT-sum BLEU ROUGE-L METEOR
-Full Model 66.29 10.95 26.97 18.80
-w/o Annotation 60.05 7.65 22.76 14.93
-w/o Phase 1 62.05 10.18 24.16 17.87
-w/o Phase 1 and Annotation 56.48 7.35 21.99 13.99
-w/o Phase 2 61.93 9.19 26.25 15.37
-w/o Phase 2 and Annotation 59.13 7.42 26.72 13.32
-TABLE VI: Performance on different levels of stripping
+Our evaluation experiments are designed to answer the following four questions:
+  1. 1.
+    RQ1 (§ VI-B). How does Malsight performance compare across two training phases and mainstream code summarization models?
+  2. 2.
+    RQ2 (§ VI-C). How do module combinations and different stripping scenarios affect Malsight’s performance, particularly the annotator module?
+  3. 3.
+    RQ3 (§ VI-D). How does the new BLEURT-sum evaluation method compare to existing code summarization evaluation metrics?
+  4. 4.
+    RQ4 (§ VI-E). How does Malsight perform when applied to real-world malicious software, and how do assessments from human reverse engineers validate its usability?
+VI-A Experimental Setup
+Our experiments run on the Ubuntu 20.04 system, equipped with one Intel Xeon Silver 4210 CPU 2.20 GHz, and two NVIDIA A40 GPUs with 125 GB RAM.
+Binary file processing tools include Radare2, IDA Pro v7.5, GCC v9.4.0, and GNU Make v4.2.1.
+Our programming language is Python v3.8.13, with transformers v4.16.2, torch v2.1.2+cu121.
+VI-B RQ1. Performance Test
+VI-B1 Code Summary Performance
+During training, the model undergoes two phases.
+The first phase involves fine-tuning malware source code (MalS) to learn malicious behavioral features.
+The second phase focuses on generating high-quality code summaries (BenignC) by incorporating semantic information from pseudocode and annotations.
+In both phases, we split the dataset into training, cross-validation, and test sets with a ratio of 7:2:1.
+We then evaluate the model at each stage to ensure its effectiveness.
+The result is shown in Table III.
+TABLE III: Score During Two Train Phases
+BLEURT-sum BLEU ROUGE-L METEOR
+Phase1 74.17 18.06 38.56 24.69
+Phase2 72.74 22.61 41.19 25.37
+CP-BCS(Baseline) 17.78 21.50 16.89 11.92
+TABLE IV: Performance comparisons on various levels of stripping
 Stripping Levels BLEURT-sum BLEU ROUGE-L METEOR AVG time (function) AVG summary length BLEURT-sum variance
@@
 * AVG summary length is measured in words.
-We designed five experiments to test different combinations for the code summarization task, as follows: (1) Removing the annotation module and summarizing without any annotation, labeled as “w/o Annotation”; (2) Canceling the phase 1, labeled as “w/o Phase ”; (3) Canceling the phase 1 and removing the annotation module, labeled as “w/o Phase 1 and Annotatio”; (4) Canceling the phase 2, labeled as “w/o Phase 2”; (5) Canceling the phase 2 and removing the annotation module, labeled as “w/o Phase 2 and Annotation”.
+Since each phase uses a different dataset, their effectiveness should be verified independently rather than by direct score comparison.
+Considering there is no universally accepted threshold for model evaluation scores, we provide the performance of another state-of-the-art binary code summarization model (CP-BCS [55]) as a baseline.
+CP-BCS is tested on its own dataset, which consists of binaries compiled with GCC 7.3.0 for x86 architecture (32-bit) and then stripped.
+Our results in both phases demonstrate higher usability compared to one of the existing method CP-BCS.
+Also, during the training process of transfer learning, we try different training data proportions and select the optimal proportion, as shown in Table V.
+TABLE V: Performance in different training data proportion
+Proportion BLEURT-sum BLEU ROUGE-L METEOR
+1:1 59.12 8.98 23.04 16.14
+1:2 59.81 9.15 23.50 16.42
+3:4 59.53 8.68 22.95 16.39
+4:3 62.14 9.80 25.11 16.88
+* Limited by MalS, proportion 2:1 is unavailable.
+* AVG summary length is measured in words.
+After completing the training step, we use MalP as the test set (it is the only pseudocode summary dataset built for malware that we know until now).
+We compare Malsight to existing pseudocode summarization methods and prompt-based general-purpose large model on dataset MalP.
+Simultaneously, we built three versions of data for MalP: not-stripped, demi-stripped (only stripping the function name), and all-stripped (stripping all identifiers inside the function).
+The closest version of the dataset to the real world, which is the results of the all-stripped version are shown in Table II.
+We conduct experiments using BLEURT-sum, BLEU, ROUGE, and METEOR (the usability of BLEURT-sum is substantiated in the referenced paper) to carry out experiments on Malsight, WizardLM [56], and Code Llama [57], etc.
+We also measure model performance using criteria such as summary length, variance, and processing time, as these factors can directly impact algorithm-based evaluations.
+(a) Malsight vs Code Llama-7b
+(b) Malsight vs WizardCoder-15b
+Figure 6: Data distribution. The segmentation distribution yielded by the proposed Malsight is more stable than the ones of the previous code-Llama.
+General-purpose large models, trained on extensive code corpora, offer an advantage in code summarization tasks, with Code-Llama performing the best, achieving a BLEURT-sum score of 56.00.
+Malsight’s innovative fine-tuning approach offers a significant advantage in summarizing malware code, with its annotation generation effectively bridging the knowledge gap between specialized and general-purpose models.
+It has been revealed that Malsight achieved the highest scores among all evaluation indexes.
+This conclusion can be corroborated by Fig.
+6.
+The figure shows that Malsight has better score evaluation results and less variance, i.e., more stable code summary output, on test results stripped of data.
+VI-B2 Annotation Extraction Effect
+The annotation extraction model is trained and tested on the AnnoS dataset, shown in Table VI.
+The annotated extraction model is trained and tested on the AnnoS dataset to complete the sequence SL task, dividing the data into normal codes (represented by N-label), important function APIs (represented by A-label), and important strings (represented by S-label).
+The model test produces the confusion matrix shown in Table VI below.
+TABLE VI: Annotation extraction confusion matrix
+Actual/Predicted N-label(Predicted) A-label(Predicted) I-label(Predicted)
+N-label(Actual) 1,847,647 9,345 11,721
+A-label(Actual) 15,865 116,099 1,784
+I-label(Actual) 21,639 1,688 38,933
+The results show that the accuracy of the model is 96.99%, which is comparable to the results obtained by manual annotation.
+Inevitably, some functions have ambiguous function names, or strings that appear to have certain characteristics, leading to negligible incorrect labeling.
+VI-C RQ2. Ablation Study
+To verify the usability of each module in Malsight, we conduct ablation experiments on the annotation module and the two phases of code summarization.
+Further, We use MalP as a test dataset, experimenting with different module combinations.
+TABLE VII: Ablation experiment
+Model BLEURT-sum BLEU ROUGE-L METEOR
+Full Model 66.29 10.95 26.97 18.80
+w/o Annotation 60.05 7.65 22.76 14.93
+w/o Phase 1 62.05 10.18 24.16 17.87
+w/o Phase 1 and Annotation 56.48 7.35 21.99 13.99
+w/o Phase 2 61.93 9.19 26.25 15.37
+w/o Phase 2 and Annotation 59.13 7.42 26.72 13.32
+VI-C1 Module Ablation
+As shown in Table VII, the modules in Malsight complement each other effectively.
+This ablation study is conducted on the not-stripped version of MalP to minimize result fluctuations due to the absence of dynamic annotation.
+We conduct experiments under five settings: (1) Removing the annotation module and summarizing without any annotation, labelled as “w/o Annotation”; (2) Canceling phase 1, labelled as “w/o Phase ”; (3) Canceling phase 1 and removing the annotation module, labelled as “w/o Phase 1 and Annotation”; (4) Canceling the phase 2, labelled as “w/o Phase 2”; (5) Canceling the phase 2 and removing the annotation module, labelled as “w/o Phase 2 and Annotation”.
 The results indicate that different degrees of ablation have varying negative impacts on systematic performance.
@@
 Notably, the absence of the annotation generation module (static annotation generation module) significantly affects performance, demonstrating its effectiveness.
-6.3.2 Annotater vs. Stripping
-The influence of different levels of strip on the annotation module continues to be explored, and the results are shown in Table VI.
+VI-C2 Annotater vs. Stripping
+The influence of different levels of the strip on the annotation module continues to be explored, and the results are shown in Table IV.
 Experiments have shown that a higher degree of stripping creates a tolerable performance degradation in the absence of annotation, demonstrating the robustness of Malsight.
-After the annotation module was ablated, it produced a 9.41% performance degradation (measured by a BLEURT-sum score) in the not-stripped test and nearly 12% in the all-stripped test, which proved that the annotator effectively resisted the adverse conditions caused by stripping.
-Figure 7: Annotation Ablation. The Annotation generation module adds more semantic information to sentence A, making it able to output function functionality smoothly, but after ablation, it cannot.
-In the generated results, we selected relatively representative sentences, as shown in Figure 7.
+After the annotation module is ablated, it produces a 9.41% performance degradation (measured by a BLEURT-sum score) in the not-stripped test and nearly 12% in the all-stripped test, which has proved that the annotator effectively resists the adverse conditions caused by stripping.
+Figure 7: Annotation ablation. The annotation generation module enriches sentence A with semantic information, enabling it to convey function functionality, which is lost after ablation.
+In the generated results, we select relatively representative sentences, as shown in Fig.
+7.
 In the yellow box, the output from the generation module without annotations shows that the large model is highly susceptible to biases due to hallucination [58].
@@
 To mitigate this issue, we incorporate annotations, which not only enhance the model’s ability to summarize behavior, structure, and application scenarios but also reduce the nonsensical outputs caused by these illusions.
-6.4 RQ3. Evaluation Algorithm Test
+VI-D RQ3. Evaluation Algorithm Test
+As previously mentioned, we use positive and negative samples to test the effectiveness of these evaluation methods by assessing their ability to distinguish between the two.
+The performance of existing methods, illustrated in Fig.
+8, shows varying degrees of crossover between positive and negative samples for BLEU, METEOR, Word2vec, and MoverScore.
+These four methods represent the current approaches in words’ overlap measure(BLEU and METEOR) and words’ embedding measure(word2vec and MoverScore), respectively.
 (a) BLEU
@@
 (d) MoverScore
-Figure 8: Relevant method test results.
+Figure 8: Test results of popular text similarity algorithms.
 The evaluation effect of two common algorithms and ML-based evaluation methods is not ideal, and their positive samples and negative samples show a certain degree of crossover.
-The EvaS dataset was constructed to evaluate our code summary evaluation method alongside other popular methods.
-As previously mentioned, we use both positive and negative samples to test the effectiveness of these evaluation methods by assessing their ability to distinguish between the two.
-The performance of existing methods, illustrated in Figure 8, shows varying degrees of crossover between positive and negative samples for BLEU, METEOR, word2vec, and MoverScore.
-These four methods represent the current approaches in words’ overlap measure(BLEU and METEOR) and words’ embedding measure(word2vec and MoverScore), respectively.
-In Figure 8, we have established two types of decision boundaries: one being orthogonal to the X-axis (i.e., distinguishing between positive and negative samples by setting a threshold), and the other being a linear function forming a slanted line.
-The same process is represented by BLEURT-sum as shown in Figure 9, which shows the superiority of our method.
-Figure 9: BLEURT-sum performance. Compared with other methods, BLEURT-sum has significantly better ability to distinguish between positive and negative samples.
-Specifically, we calculated an F1-score for these six methods (ROUGE-L is not shown in the Figure 8) on the positive and negative sample classification task of code summary sentences to measure their ability.
-Our method achieved an F1-score exceeding 0.9999, significantly outperforming all other evaluation methods.
-Among the existing methods, METEOR performed the best with an F1-score of 0.9811, while BLEU had the lowest performance at only 0.85 as the lowest.
-None of the other methods achieved an F1-score above 0.95.
-Notably, our dataset was not meticulously curated; the negative samples were entirely random sentences with different meanings, which differ from real-world scenarios where the differences in meaning might be subtler.
-An F1-score below 95% may therefore indicate an unacceptable tendency to misclassify.
-This is because a 95% F1-score means that approximately 1 in 10 samples are incorrectly judged while upholding a high recall rate.
+In Fig.
+8(a), we have established two types of decision boundaries: one being orthogonal to the X-axis (i.e., distinguishing between positive and negative samples by setting a threshold), and the other being a linear function forming a slanted line.
+The same process is represented by BLEURT-sum as shown in Fig.
+9, which shows the superiority of our method.
+Specifically, we calculate the F1-score for these six methods (ROUGE-L is not shown in Fig.
+8) on the positive and negative sample classification task of code summary sentences to measure their ability.
+Our method achieves an F1-score exceeding 0.9999, significantly outperforming all other evaluation methods.
+Among the existing methods, METEOR performs best with an F1-score of 0.9811, while BLEU gets the lowest performance at only 0.85.
+None of the other techniques achieve an F1-score above 0.95.
+(a) BLEURT-sum
+(b) ROC Curve (Scaled)
+Figure 9: BLEURT-sum performance.
+BLEURT-sum has a significantly better ability to distinguish between positive and negative samples(a).
+Simultaneously, the ROC curve demonstrates a commendable AUC.(b)
+Notably, our dataset is not meticulously curated, the negative samples are entirely random sentences with different meanings, which differ from real-world scenarios where the differences in meaning might be subtler.
+An F1-score below 95% indicates a tendency to misclassify, with approximately 1 in 10 samples judged incorrectly.
 As an evaluation method, the misjudgment can be further amplified by the model results evaluated using the evaluation scheme in the following works.
-6.5 RQ4. Real World Experiments
-6.5.1 Human Evaluation Experiments
-So far, we have evaluated Malsight’s performance at the function level.
-However, this approach did not allow us to fully showcase the capabilities of the dynamic annotation module.
-Additionally, real-world malware predominantly manifests in executable files.
-Hence, we manually analyzed 10 real-world malware samples and selected 79 critical functions out of them, three experienced reverse engineers add summaries to these functions based on discussion.
-The complete workflow of Malsight was applied to these functions and the outputs was evaluated using both BLEURT-sum and human evaluation metrics.
-We further calculated the variance deviation of all calculated results as a reference to demonstrate the stability of our solution and the time required for code summarization for a single malware sample.
-For the human evaluation metrics, we invited ten evaluators, including five reverse engineers who are rich in experience, three experienced, and two beginners.
-They were asked to focus on evaluating the usability of the code summarization results and provide a score of 0, 0.5 or 1, indicating whether the summary corresponded to the original code.
-It is noteworthy that due to the subjectivity of human evaluation of the summaries, different evaluators may have varying opinions on the summary of the same code.
-Therefore, we opted to evaluate usability only and did not solicit more detailed scores beyond 0,0.5 and 1, respectively corrrepond to usable, partially usable and unusable.
-6.5.2 Performance From Different Evaluators
-After three sets of scores, we obtained the human assessment scores shown in Table VII.
-Experienced evaluators tend to give more conservative and stable scores, while less experienced evaluators give a wider range of scores with higher variance.
-Further, we take the arithmetic average of the scores received by these evaluators as the final human assessment score.
-TABLE VII: Evaluation of Different Categories of Evaluators
+VI-E RQ4. Real World Experiments
+VI-E1 Human Evaluation Experiments
+So far, we have evaluated Malsight’s performance at the function level, but this hasn’t fully demonstrated the dynamic annotation module’s capabilities.
+Since real-world malware predominantly appears as executables, we manually analyzed ten malware samples and selected 79 critical functions.
+Three experienced reverse engineers collaboratively provided summaries for these functions.
+Malsight’s complete workflow is applied to these samples, with outputs evaluated using both BLEURT-sum and human evaluation metrics.
+We also calculate the variance deviation to assess the stability of the solution and measure the time required for summarizing a single malware sample.
+For human evaluation, we invite ten evaluators, including five expert reverse engineers, three experienced engineers, and two beginners.
+They rate the summaries based on usability with scores of 0, 0.5, or 1, corresponding to unusable, partially usable, and usable, respectively.
+Given the subjectivity of human evaluation, we focus solely on usability rather than providing more detailed scores.
+VI-E2 Performance From Different Evaluators
+After three sets of scores, we obtain the human assessment scores shown in Table VIII.
+Experienced evaluators tend to give more conservative and stable scores, while less experienced evaluators provide a wider range of scores with higher variance.
+Further, we take the arithmetic average of the scores from the evaluators as the final human assessment score.
+TABLE VIII: Results of Different Evaluators
 Rich in experience Experienced Beginner
@@
 Variance 173.12 197.49 237.44
-Table b shows the mean value of BLEURT-sum and the mean value of human evaluation on 79 labeled functions.
-The human evaluation score of 59.87 means that most functions have exceeded 0.5, i.e.
-the partially usable standard.
-TABLE VIII: Real Malware Performance
-BLEURT-sum Human Assessment Score AVG time (file)
+Table IX shows the mean value of BLEURT-sum and the mean value of human evaluation on 79 labeled functions.
+The human evaluation score of 59.87 means that most functions have exceeded the partially usable standard.
+TABLE IX: Real Malware Performance
+BLEURT-sum Human Evaluation Score AVG time (file)
 Score 47.22 59.87 1.90
-Variance 161.61 0.06 1.21
+Variance 161.61 193.34 1.21
 * AVG time (per file) is measured in hours.
-Figure 10: Real-world Performance. While the ratio of human assessment score to BLEURT-sum remained positive, Malsight is able to obtain higher assessment results.
-Further, we formed a data point by combining a BLEURT-sum score with a human evaluation score for the same function’s code summary, thus, we plotted a distribution of 79 data points in Figure 10.
-It indicates that BLEURT-sum has a positive linear correlation with human evaluation metrics, which shows the rationality of BLEURT-sum.
-Furthermore, the data distribution indicates that, across the two distinct evaluation metrics, the majority of the evaluation scores for Malsight’s outputs are concentrated within the middle to high range, which shows that our scheme is also usable in the real world.
-7 Case Study
-In section RQ4, we evaluate Malsight’s performance on 10 real-world malware.
-To provide more direct evidence and analyze the reasons why Malsight behaves differently on different functions, three functions are presented in this section to discuss the process that Malsight goes through behind the output that is usable, the output that is partially usable, and the output that is unusable.
-Figure 11: Redacted pseudocode in Case 1.
+(a) Scatter Plot
+(b) Violin Plot
+Figure 10: Real-world Performance. While the ratio of human assessment score to BLEURT-sum remained positive, Malsight can obtain higher assessment results.
+Further, we form a data point by combining a BLEURT-sum score with a human evaluation score for the same function’s code summary.
+Thus, we plot a distribution of 79 data points in Fig.
+10(a).
+It indicates that the BLEURT-sum has a positive linear correlation with human evaluation metrics, which shows its rationality.
+Furthermore, the data distribution in Fig.
+10(b) indicates that, across the two distinct evaluation metrics, the majority of the evaluation scores for Malsight’s outputs are concentrated within the middle to high range, which shows that our scheme is also usable in the real world.
+VII Case Study
+In Section RQ4, we evaluate Malsight’s performance on ten real-world malware samples.
+To provide deeper insights into Malsight’s varying performance across functions, we present two functions to illustrate the process behind generating a high-quality output and a lower-quality one.
+Figure 11: Pseudocode of case 1&2. The two real malware functions produced varying results, with the model’s performance slightly weaker on the longer function due to length limitations.
 Case 1: sub_401021.
@@
   • Human Summary: Escalating the privileges of the virus process.
-Malsight’s summary of the function shown in Figure 11 is considered usable.
+Malsight’s summary shown in Fig.
+11(right) is considered high-quality.
 In this case, Malsight summary accurately identifies “setting privileges” as the key point from a lengthy function, which aligns with the task of “privilege escalation” in the human summary.
-Figure 12: Redacted pseudocode in Case 2.
 Case 2: sub_401D30.
@@
   • Human Summary: Connect to a malicious domain and make a request to retrieve data using HTTP.
-Malsight’s summary of the function shown in Figure 12 is considered partially usable.
-Malsight summary identifies that the function uses an HTTP connection to a specific server and operates on the data, which is consistent with the human summary.
-However, due to the lack of some background information, it fails to recognize that the specific server is malicious.
-Figure 13: Redacted pseudocode in Case 3.
-Case 3: TimerFunc.
-  • Malsight Summary: Get the command line, the code defines a function that handles or processes data related to a hexadecimal address, retrieves information about Redis commands and modules, Call the NtTib.
-ExceptionList, the code defines a function that handles or processes data related to a hexadecimal address, retrieves information about Copy the setup.exe, the code defines a function that handles or processes data related to a hexadecimal address, retrieves information about Redis commands and modules Setup the setup, the code defines a function that handles or processes data related to a hexadecimal address, retrieves information about Redis commands and modules.
-  • Human Summary: Checking if the autorun.ini file exists; if not, create it and configure it to auto-run setup.exe as specified in the configuration file.
-Malsight’s summary of the function shown in Figure 13 is considered unusable.
-The function is very long, contains numerous function calls, and includes many redundant functions that hinder analysis, resulting in a poor Malsight summary.
-8 Discussion
-8.1 Ethics
-The construction of the dataset MalS and MalP complies with GitHub’s open-source license agreements.
-We only used open-source repositories with GitHub open-source licenses and downloaded these open-source codes within the rate limit set by GitHub to minimize interference with GitHub’s servers to the lowest extent possible.
-Additionally, We enlisted volunteers to review the MalS and MalP datasets, as well as the 10 malware samples mentioned in RQ4.
-With the volunteers’ consent, we adopted their review results as the final dataset.
-8.2 Limitations
+Malsight’s summary of the function shown in Fig.
+11(left) is considered lower-quality.
+The summary correctly identifies the function’s use of an HTTP connection to a server and data operations, aligning with the human summary.
+However, it misses that the server is malicious due to lacking background information.
+Additionally, the model’s output tends to be lengthy and repetitive due to the function’s large code body.
+We believe this is related to the model’s parameter size and the input length it can process, using a model with a larger parameter count could further improve Malsight’s performance.
+VIII Discussion
 In this section, we delve into practical issues based on our analysis of the experimental results. Additionally, we explore aspects not covered in our work and propose potential solutions.
 Real-World Malware vs.
-Malware Function Summaries: In our exploration of real-world malware, the BLEURT-sum evaluation score of the code summary was approximately 30% lower than the experimental score obtained by MalP in Section 6.2.
-The reason could be that our code summarization model lacks the ability to summarize longer functions which occasionally appear in real-world malware.
-Our analysis shows that about 15% of the functions in the malware contain more than 1000 tokens (while the majority of the functions in the training set stay below 300 tokens), which is likely due to different optimization configurations during compilation.
-These lengthy function bodies introduce a lot of information and noise, which makes it difficult for the model to extract and summarize the critical code fragments.
-Therefore, to better align with real malicious code summaries, researchers should consider doing code summarization work in units of code fragments instead of functions.
-However, identifying and summarizing important, human-interpretable segments in them requires large amounts of labeled data, or compositional functional insights from dynamic debugging, which poses significant challenges.
+Function Summaries: In our evaluation of real-world malware, the BLEURT-sum score for code summaries was approximately 30% lower than the experimental score from MalP in Section VI-B.
+This discrepancy may arise from the model’s difficulty in summarizing longer functions, which are more common in real malware.
+Our analysis reveals that 15.8% of the malware functions exceed 1000 tokens, while most training set functions are under 300 tokens.
+This length introduces excessive information and noise, complicating the extraction of critical code fragments.
+Therefore, we investigate code summaries at the code fragment level, discovering that the model offers enhanced insights in this format.
+We propose adopting a large model agent approach for generating summaries of code fragments, which can then be integrated to form function-level summaries.
+Furthermore, utilizing additional labelled data from manual dynamic debugging can further enhance the results.
 Summary for assembly language rather than pseudocode: Efforts have been made to recover information from assembly code and to generate function summaries [59].
@@
 Thus, starting directly from assembly language is considered a viable solution.
-In our comparison of assembly code and pseudocode representations, we found that assembly code is challenging for models pre-trained on high-level languages to understand [32].
+In our comparison of assembly code and pseudocode representations, we find that assembly code is challenging for models pre-trained on high-level languages to understand [32].
 The structural features of assembly code are completely different from those of high-level languages.
 Fine-tuning a model using pseudocode can confuse the model’s understanding of function-level structural features, leading to unacceptable error output, possibly due to insufficient datasets.
-We train an embedding model specifically for assembly language and build a large dataset at the assembly language level for malware.
-This method has the potential to outperform general-purpose large models, which typically have a poor understanding of assembly code.
-9 Conclusion
-In this paper, we introduced Malsight, a novel framework for binary malware summarization by exploring malicious source code and benign pseudocode.
-The proposed Malsight involves two datasets MalS and MalP, an LLM-based summary model, and an evaluation metrics.
-Experimental results on three datasets show the effectiveness of the proposed framework.
-Future work includes the application of the proposed framework to more downstream tasks.
+On the other hand, the presence of lengthy assembly code can significantly increase the inference overhead for code summarization models.
+Moreover, the substantial differences in assembly code produced by various compilers may introduce data shifts.
+This issue is mitigated by consistently treating all code as pseudocode during processing.
+IX Conclusion
+In this paper, we addressed the challenges of binary malware summarization by introducing Malsight, a novel framework that leverages insights from both malicious source code and benign pseudocode.
+Malsight was designed to overcome the limitations of existing methods in handling entangled logic and stripped semantics commonly found in malware binaries.
+Our contributions include the development of two key datasets, MalS and MalP, specifically constructed to facilitate malware summarization.
+We also introduced MalT5, a lightweight LLM-based model tailored for this task, and proposed BLEURT-sum, an innovative evaluation metric that enhances code summarization performance beyond traditional word overlap metrics.
+Experimental results on both test sets and real-world malware datasets demonstrated the effectiveness of Malsight compared to other state-of-the-art methods.
+The framework successfully addressed critical issues such as recovering meaningful summaries from complex and stripped pseudocode, underscoring its robustness in handling real-world challenges.
+Future research may focus on improving the model’s ability to generate concise summaries for longer code sequences and applying the Malsight framework to more complex downstream tasks.
+Additionally, extending the approach to further real-world malware scenarios remains a crucial direction for continued development.
 Appendix A Analysis of Traditional Similarity Evaluation Metric
-The criteria for evaluating sentence similarity should focus on two key aspects: similar sentences should correspond to higher scores, while dissimilar sentences should correspond to lower scores.
-Below, we introduce some of the shortcomings of BLEU and other traditional metrics (ROUGE, METEOR) in these aspects.
-A.1 Limitations of Overlap Based Methods
 Problems caused by BLEU algorithm: The formula BLEU uses to calculate sentence similarity is shown in equation 20:
@@
 Subjectively, when two target sentences have a higher degree of overlap in a specific word, they are given a higher ROUGE score.
-METEOR designs on the basis of rouge, following equation 20, 21 and 22.
+METEOR designs on the basis of rouge, following equation 23, 24 and 25.
 $F_{mean}=\frac{(1+\beta^{2})PR}{R+\beta P}$ (23)
@@
 Interestingly, we found that expanding the stopword list appropriately can enhance the performance of word overlap-based methods, especially METEOR.
-A.2 Limitations of Embedding Based Methods
-In recent years, using word embedding-based methods such as word2vec to determine sentence similarity has become quite popular.
-For instance, the pseudocode for the word2vec process is illustrated in Algorithm 1.
-Algorithm 1 word2vec
-1:The reference sentence $ref$ and the candidate sentence $cad$
-2:the similarity $score$ of $ref$ and $cad$
-3:function Main($cad,ref$)
-4:     cad.vector = getVector($cad$)
-5:     ref.vector = getVector($ref$)
-6:     $\text{score}=\frac{\text{cad}\cdot\text{ref}}{||\text{cad}||\cdot||\text{ref}||}$
-7:     return score
-8:end function
-9:function getVector($sentence$)
-10:     words = tokenize(sentence)
-11:     sentence.vector = [0,0,..]
-[…]
-13:         sentence.vector += word.vector
-14:     end for
-15:     return sentence.vector
-16:end function
-When calculating the similarity between two sentences, word2vec adds the vectors of each word based on the tokenized results and uses the summed vector as the sentence representation.
-Finally, cosine similarity is employed to determine the similarity between the sentences.
-A direct drawback of this approach is that it loses all word order information and heavily depends on the accuracy of the embedding algorithm.
-Appendix B Reverse CFG Sorting Algorithm
-Algorithm 2 REsort
-1:The CFG graph $G_{M}$ for a malware binary file.
-2:The reverse topsort list $L_{G}$ of the function call graph.
-3:function REsort($G_{M}$)
-4:     for each vertex $v$ in $G_{M}.vertices$ do
-5:         if $!v.seen$ then
-6:              new $G_{M}^{\prime}$
-7:              $G_{M}^{\prime}.vertices\leftarrow$ call Tarjan($v$, $G$)
-8:              //Tarjan() traverses all strongly connected components in $G_{M}$ and contracts vertices as $G_{M}^{\prime}.vertices$.
-9:         end if
-10:     end for
-11:     $G_{M}^{\prime}\leftarrow$ call BuildTarGraph($G_{M}^{\prime}.vertices$, $G_{M}$)
-12:     //BuildTarGraph() build a new directed acyclic graph $G_{M}^{\prime}$ using $G_{M}^{\prime}.vertices$ and $G_{M}.edges$
-13:     $L_{G_{M}^{\prime}}\leftarrow$ call RetopSort($G_{M}^{\prime}$)
-14:     //RetopSort() obtains the reverse topological sorting sequence of $G_{M}^{\prime}$
-15:     $dist[]\leftarrow$ call Dijkstra($G_{M}$)
-16:     //Dijkstra() computes the multi-source shortest paths in graph $G_{M}$ from vertices with an in-degree of 0 to each vertix.
-17:     $L_{G_{M}}\leftarrow[]$
-18:     for each vertex $v^{\prime}$ in $L_{G_{M}^{\prime}}$ do
-19:         $mindist\leftarrow\infty$, $idx\leftarrow-1$
-20:         //$idx$ determines the starting vertex for DFS in each strongly connected component in $G_{M}$
-21:         for each $ver$ in $v^{\prime}.subvertex$ do
-22:              if $dist[ver]<mindist$ then
-23:                  $mindist\leftarrow dist[ver]$, $idx\leftarrow ver$
-24:              end if
-25:         end for
-26:         $L_{G_{M}}$.append(DFS($idx$, $v^{\prime}$))
-27:         //DFS() derives a traversal order in each strongly connected component in $G_{M}$ as part of the final reverse topological sorting.
-28:     end for
-29:end function
-TABLE IX: Performance in different training data proportion
+TABLE X: Performance in different training data proportion
 Proportion BLEURT-sum BLEU ROUGE-L METEOR AVG summary length BLEURT-sum variance
@@
 * AVG summary length is measured in words.
-Algorithm 2 presents our function topological sorting algorithm, with the objective of constructing the reverse topological sorting of the function call graph $G_{M}$.
-Due to the presence of multi-cycles (i.e., strongly connected components) in $G_{M}$, it is not feasible to perform reverse topological sorting directly.
-To address this, we utilize the Tarjan algorithm to contract $G_{M}$ into a new directed acyclic graph $G_{M}^{\prime}$, where reverse topological sorting can be applied to obtain an initial function traversal list $L_{G_{M}^{\prime}}$, with each node representing a strongly connected component in $G_{M}$.
-Further, to achieve topological sorting for $G_{M}$, we employ the DFS algorithm to sort each node within the strongly connected components.
-Since our aim is to reverse-sort the entire graph $G_{M}$, with functions closer to the start function preferred towards the end of $L_{G_{M}}$, we utilize the Dijkstra algorithm to compute the multi-source shortest paths $dist[]$ from all nodes with zero in-degree to other nodes (referred to as ’multi-source’ because the function call graph $G_{M}$ may not necessarily be connected).
-Subsequently, by traversing $L_{G_{M}^{\prime}}$ sequentially, for each strongly connected component, the node with the minimum $dist$ value is selected as the initial node $idx$ for DFS.
-Leveraging the inherent stack property of DFS, $idx$ ensures its placement at the end of $L_{G_{M}}$.
-Finally, $L_{G_{M}}$ represents the desired function traversal order.
-Appendix C Performance in different proportion of two-phase fine-tuning training data size
+TABLE XI: Summarization examples on TLS-related Go code by various LLMs.
+Method (LLM) Summary
+BinT5 Go TLS connection handling with runtime checks.
+HexT5 TLS handshake timeout configuration.
+WizardCoder-15B This Go code sets up TLS handshake timeouts for HTTP servers, and it ensures secure connections by configuring read/write deadlines during the TLS handshake phase, preventing the server from hanging indefinitely.
+Additionally, it includes runtime checks to ensure compatibility.
+Code Llama-7B The code enforces TLS handshake timeouts to prevent hanging connections, setting deadlines during negotiation.
+It includes runtime checks to ensure compatibility while keeping encrypted chats from dozing off mid-conversation.
+CodeT5+ This code sets TLS timeouts for HTTP servers, ensuring security.
+deepseek-llm-7b-chat This code formats an integer, performs string operations, initializes TLS connection, handles time, and conditionally closes a connection or returns a Go interface.
+Appendix B Performance in different proportion of two-phase fine-tuning training data size
 We sampled a total of 140,000 pieces of data in different proportions from the two datasets (MalS and BenignC) for two phases of fine-tuning.
-The results of the evaluation of model performance in each training data proportion are shown in the Table IX.
+The results of the evaluation of model performance in each training data proportion are shown in the Table X.
 In the experimental part of the text, we choose the ratio of 4:3 with the best effect as the actual training set ratio of Malsight.
+Therefore, in the subsequent experiments, we adopt this configuration as the default training set composition for Malsight.
+Figure 13: Application of MALSIGHT to other binary code summarization tasks.
+Appendix C Case Study Supplement
+C-A Malsight for other binary code summarization task
+We complement this with a case study on a pseudo-code summarization task on binary files written and compiled in Go (Golang) (shown in Fig. 13), aiming to demonstrate the generalizability of Malsight.
+Case: Go Binary Psuedo-code.
+  • Malsight Summary: Initiate TLS connection, the code formats an integer into a string, performs various string operations including concatenation, initializes a TLS connection, handles time-related operations such as adding time, and conditionally closes the connection or returns a handle based on runtime checks.
+C-B Application of other LLMs to the binary summarzation task
+We would like to show the application of other large language models to the binary summarization task.
+We will use the same pseudo-code snippet (Fig.
+13) shown in the previous subsection.
+The results are presented in Table XI.